### v4

# Measuring Evidence against Exchangeability and Group Invariance with E-values

Abstract
We study e-values for quantifying evidence against exchangeability and general invariance of a random variable under a compact group. We start by characterizing such e-values, and explaining how they nest traditional group invariance tests as a special case. We show they can be easily designed for an arbitrary test statistic, and computed through Monte Carlo sampling. We prove a result that characterizes optimal e-values for group invariance against optimality targets that satisfy a mild orbit-wise decomposition property. We apply this to design expected-utility-optimal e-values for group invariance, which include both Neyman–Pearson-optimal tests and log-optimal e-values. Moreover, we generalize the notion of rank- and sign-based testing to compact groups, by using a representative inversion kernel. In addition, we characterize e-processes for group invariance for arbitrary filtrations, and provide tools to construct them. We also describe test martingales under a natural filtration, which are simpler to construct. Peeking beyond compact groups, we encounter e-values and e-processes based on ergodic theorems. These nest e-processes based on de Finetti’s theorem for testing exchangeability.
Keywords: permutation test, group invariance test, anytime valid inference, post-hoc valid inference, $e$-values, sequential testing.
1 Introduction
Testing group invariance is an old and fundamental problem in hypothesis testing. It covers many non-parametric tests, including permutation tests, conformal inference, various popular multiple testing methods, many causal inference methods, and even the $t$-test. Tests for group invariance are attractive and widely used, as invariance properties under a null hypothesis are often easy to defend. Rejecting this invariance then also rejects the null hypothesis of interest.
Up to the present, only a limited class of group invariance tests has been explored, of a form inspired by the traditional Neyman–Pearson framework of testing. The primary contribution of this manuscript is to go beyond this traditional framework by measuring evidence against group invariance with e-values (Shafer, 2021; Vovk and Wang, 2021; Howard et al., 2021; Ramdas et al., 2023b; Grünwald et al., 2024; Koning, 2024c).
Before we detail our contributions, we first briefly discuss the traditional permutation test, which serves as a prototypical example of a traditional group invariance test. We follow this with a short primer on e-values.
1.1 Traditional permutation test
Consider the random variable $X^{n}=(X_{1},\dots,X_{n})$, and suppose we are interested in testing whether it is exchangeable:
$H_{0}:X^{n}\text{ is exchangeable}.$
Here, exchangeability means that $X^{n}$ is equal in distribution to every permutation $PX^{n}$ of its elements. As an example, $X^{n}$ is exchangeable if its components $X_{1},\dots,X_{n}$ are i.i.d., but there also exist non-i.i.d. exchangeable distributions.
Given some test statistic $T$, a traditional ‘permutation p-value’ is given by
$p(X^{n})=\mathbb{P}_{\overline{P}_{n}}\left(T(\overline{P}_{n}X^{n})>T(X^{n})\right),$
where $\overline{P}_{n}\sim\textnormal{Unif}(\mathfrak{P}_{n})$ is uniformly distributed on the permutations $\mathfrak{P}_{n}$ of $n$ elements.This p-value can be understood as the proportion of test statistics calculated from the rearranged (‘permuted’) data that exceed or match the original test statistic.
Instead of formulating a p-value, we can equivalently formulate a traditional permutation test
$\varepsilon_{n}^{(\alpha)}(X^{n})=\mathbb{I}\{p(X^{n})\leq\alpha\}/\alpha,$
where we follow Koning (2024c) in modeling a level $\alpha$ test $\varepsilon_{n}^{(\alpha)}$ as a map to the interval $[0,1/\alpha]$, so that $\varepsilon_{n}^{(\alpha)}(X^{n})=1/\alpha$ indicates a rejection at level $\alpha$ and $\varepsilon_{n}^{(\alpha)}(X^{n})=0$ a non-rejection.
Permutation tests and p-values are well-known to be valid in finite samples:
$\sup_{n,\alpha}\mathbb{E}^{\mathbb{P}}[\varepsilon_{n}^{(\alpha)}(X^{n})]\equiv\sup_{n,\alpha}\mathbb{E}^{\mathbb{P}}[\mathbb{I}\{p(X^{n})\leq\alpha\}/\alpha]\leq 1,$ (1)
and every exchangeable distribution $\mathbb{P}$, or, equivalently, $\mathbb{P}(p(X^{n})\leq\alpha)\leq\alpha$, for all $n$ and $\alpha$. In fact, this can even be made to hold with equality by breaking ties through randomization.
Permutation tests are a special case of more general group invariance tests, which are obtained by simply replacing the group of permutations $\mathfrak{P}_{n}$ by some other compact group $\mathcal{G}_{n}$ that acts on our sample space. In Section 2, we cover the necessary background on compact groups and their actions on sample spaces.
1.2 Primer on e-values
Traditional group invariance tests are binary by construction: such a test either rejects the hypothesis at level $\alpha$ or not: $\varepsilon_{n}^{(\alpha)}=1/\alpha$ or $\varepsilon_{n}^{(\alpha)}=0$. Recently, there has been much interest in moving away from such binary tests towards ‘fuzzy’ tests that exploit the entire domain $[0,1/\alpha]$ or even $[0,\infty]$ (Koning, 2024c). These tests have been popularized under the name e-values (Shafer, 2021; Howard et al., 2021; Grünwald et al., 2024). Such e-values are to be used as a continuous measure of evidence against the hypothesis, where its value on $[0,1/\alpha]$ or $[0,\infty]$ is directly interpreted as evidence. The introduction of the e-value has led to a series of breakthroughs in sequential testing, multiple testing and even in single-hypothesis testing.
For e-values bounded to $[0,1/\alpha]$, the Neyman–Pearson lemma tells us that binary e-values (tests) automatically arise when maximizing the power $\mathbb{E}^{\mathbb{Q}}[\varepsilon]$ for simple null hypotheses, and also when testing group invariance (Lehmann and Stein, 1949). To obtain non-binary e-values, we must consider other power-targets, such as $\mathbb{E}^{\mathbb{Q}}[\log\varepsilon]$. E-values that maximize this target have also been dubbed ‘log-optimal’, ‘GRO’ or ‘numeraire’ (Shafer, 2021; Grünwald et al., 2024; Larsson et al., 2024). While this log-target is most popular, more general power-targets such as maximizing some expected utility have also been studied (Koning, 2024c).
A sequential generalization of an e-value is an e-process $(\varepsilon_{n})_{n\geq 1}$. Such an e-process is said to be valid if (1) holds not just for data-independent $n$, but for stopping times:
$\sup_{\tau}\mathbb{E}^{\mathbb{P}}[\varepsilon_{\tau}(X^{\tau})]\leq 1,$
for every stopping $\tau$, and every group invariant distribution $\mathbb{P}$. Such e-processes relieve us from pre-specifying a number of observations, and permit us to continuously monitor the data and current evidence, and stop whenever we desire.
Another interpretation of the e-value may be found in its relationship to the p-value. Specifically, the reciprocal $\mathfrak{p}=1/\varepsilon$ of an e-value $\varepsilon$ is a special kind of p-value that is valid under a much stronger Type-I error guarantee (Koning, 2024a; Grünwald, 2024), also called the post-hoc level Type-I error:
$\sup_{n}\mathbb{E}_{X^{n}}\left[\sup_{\alpha}\mathbb{I}\{\mathfrak{p}(X^{n})\leq\alpha\}/\alpha\right]\equiv\mathbb{E}_{X^{n}}\left[1/\mathfrak{p}(X^{n})\right]\leq 1.$
This is stronger than the traditional Type-I error (1), as the supremum over $\alpha$ is now inside the expectation which means that it is also valid when using data-dependent significance levels $\alpha$. Indeed, $\alpha$ may be chosen post-hoc.
The smallest data-dependent level at which we reject is the p-value $\mathfrak{p}$ itself. As a consequence, we may truly ‘reject at level $\mathfrak{p}$’ and still retain a generalized Type-I error control for data-dependent significance levels (Koning, 2024a). This is certainly not permitted for traditionally valid p-values, which only offer a guarantee when compared to an independently specified level (1). To distinguish these p-values from traditional p-values, they are sometimes referred to as post-hoc p-values. Such a post-hoc p-value, and by extension the e-value, may therefore be viewed as a p-value that offers a generalized Type-I error guarantee, even when interpreted continuously.
1.3 First contribution: characterizing and computing e-values for group invariance
In Section 3, we study the characterization of e-values for group invariance for a compact group $\mathcal{G}_{n}$. There, we show $\varepsilon$ is a valid e-value for $\mathcal{G}_{n}$ invariance if and only if
$\mathbb{E}_{\overline{G}_{n}}[\varepsilon(\overline{G}_{n}x^{n})]\leq 1,$
where $\overline{G}_{n}\sim\textnormal{Unif}(\mathcal{G}_{n})$, for every $x^{n}$ in the sample space. That is, it needs to be valid under a uniform distribution, $\overline{G}_{n}x^{n}\sim\textnormal{Unif}(O_{x^{n}})$, on each ‘orbit’ $O_{x^{n}}=\{Gx^{n}:G\in\mathcal{G}_{n}\}$. In fact, as the data $X^{n}$ identifies the orbit in which it falls, we find that the e-value only needs to be valid for the uniform distribution $\textnormal{Unif}(O_{X^{n}})$ on the orbit $O_{X^{n}}$ in which the data $X^{n}$ lands.
We use this to show that an e-value is exactly valid if and only if the e-value is of the form
$\varepsilon_{T}(x^{n})=\frac{T(x^{n})}{\mathbb{E}_{\overline{G}_{n}}[T(\overline{G}_{n}x^{n})]},$
for some non-negative and appropriately integrable test statistic $T$. We find e-values of this form can be computed easily by replacing the denominator with a Monte Carlo average over i.i.d. samples from $\overline{G}_{n}$. Indeed, we show such a ‘Monte Carlo e-value’ is valid in expectation over the Monte Carlo draws. As a side contribution, we also derive a weaker condition under which traditional group invariance tests are valid.
1.4 Second contribution: optimal e-values for group invariance
In Section 4, we consider optimal e-values for group invariance for an abstract power-target. Under a mild assumption that the target monotonically aggregates local optimality targets on orbits, we find that an e-value is optimal for group invariance if it is locally optimal on each orbit. In fact, such e-values are optimal uniformly in any monotone aggregation of the local orbit-wise targets. We apply this idea to derive expected-utility-optimal e-values for group invariance, which optimize $\mathbb{E}^{\mathbb{Q}}[U(\varepsilon)]$ for some alternative $\mathbb{Q}$ and utility function $U$.
To present such optimal e-values, it is helpful to introduce the $\mathcal{G}_{n}$ invariant version $\overline{\mathbb{Q}}$ of the alternative $\mathbb{Q}$, which may be constructed by averaging the $\mathbb{Q}$-probability mass of each event over group-transformations of the event. We find that expected-utility optimal e-values may be expressed in terms of densities of $\mathbb{Q}$ and $\overline{\mathbb{Q}}$ with respect to some measure that dominates both (e.g. (+¯$(\mathbb{Q}+\overline{\mathbb{Q}})/2$). In case we additionally assume $\overline{\mathbb{Q}}\gg\mathbb{Q}$ and assume $U$ satisfies some regularity assumptions, then the expected-utility-optimal e-value is given by
$\varepsilon^{U}=(U^{\prime})^{-1}\left(\lambda^{*}\left\{\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}\right\}^{-1}\right),$
for some orbit-dependent constant $\lambda^{*}$. Specializing this to log-utility $U:x\mapsto\log(x)$ yields
$\varepsilon^{\textnormal{log}}=\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}},$
which may be interpreted as the ‘likelihood ratio’ between $\mathbb{Q}$ and invariance under the group $\mathcal{G}_{n}$. This simultaneously reveals that $\overline{\mathbb{Q}}$ may be viewed as the Reverse Information Projection (‘RIPr’) of $\mathbb{Q}$ onto the collection of $\mathcal{G}_{n}$ invariant probabilities under the KL-divergence (Grünwald et al., 2024; Lardy et al., 2024; Larsson et al., 2024). Moreover, for the utility function $U:x\mapsto x\wedge 1/\alpha$, which yields the classical Neyman–Pearson notion of ‘power’, an expected utility optimal e-value is given by
$\varepsilon_{\alpha}^{\textnormal{NP}}=\frac{1}{\alpha}\mathbb{I}\left\{\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}>c_{\alpha}\right\}+\frac{k}{\alpha}\mathbb{I}\left\{\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}=c_{\alpha}\right\},$
where $c_{\alpha}$ and $k$ are certain orbit-dependent constants. This recovers the Neyman–Pearson optimal test for group invariance derived by Lehmann and Stein (1949).
Instead of specifying an alternative on the original sample space, we also consider specifying an alternative on each orbit, as well as conditional on the data-orbit. Finally, we also consider specifying an alternative on the group, by passing the data through a unique representative inversion function, which maps it to the group. We show that this nests traditional rank and sign-based tests, by using the fact that the ranks and signs are in bijection with the group of permutations and sign-flips.
1.5 Third contribution: e-processes and test martingales for group invariance
In Section 5, we study e-processes for group invariance. We generalize the characterization of e-values to e-processes for arbitrary filtrations, and show how they may be constructed by tracking an infimum over martingales for uniform distributions $\textnormal{Unif}(O)$ on orbits $O$. We link this to our derivation of optimal e-values, by showing how optimal e-values may be used to induce such orbit-wise martingales. Moreover, we identify a key challenge when constructing e-processes: the orbit in which the data lands is not necessarily measurable at the start of the filtration, and is possibly not even measurable for any sigma-algebra in the filtration. This explains why we cannot reduce the sequential problem to testing uniformity on the orbit in which the data lands: we must instead keep track of all orbits in which the data may feasibly land.
In addition, we peek beyond compact groups by relying on an ergodic theorem for possibly non-compact groups. Here, we find e-values and e-processes may be characterized by an infimum over e-values for ergodic measures, which replace the role of uniform distributions on orbits in the compact setting. An example of an ergodic theorem is de Finetti’s theorem, where these ergodic measures are the i.i.d. probabilities, which was explored in the context of e-processes for binary and $d$-ary data by (Ramdas et al., 2022b).
In Section 6, we continue by zooming in towards a particular setup which may be viewed as sequentially testing invariance, since we consider testing whether the sequence $(X^{n})_{n\geq 1}$ is invariant under a sequence of groups $(\mathcal{G}_{n})_{n\geq 1}$. We study test martingales for this setup, which are special e-processes that satisfy a stronger condition. We show how they may easily be constructed by tailoring conditional e-values for their increments based on subgroups that stabilize the previous data.
In Appendix B, we study the impoverishment of filtrations, which is the process of deliberately moving to a less-informative filtration, usually in order to design a more powerful e-process by shrinking the set of stopping times under which it must be valid. In practice, this means we restrict ourselves to looking at some statistic of the data, instead of at the underlying data itself. We show how we may generally impoverish filtrations in the context of group invariance, by using a statistic that is equivariant under a subgroup of the group. The problem then reduces to testing invariance of the statistic under the subgroup.
We then focus on equivariant statistics for which the subgroup only has a single orbit on its codomain. We find that an example of such an equivariant statistic is the unique representative inversion that maps the data to the group. This nests the idea studied by Vovk (2021) to pass to the ranks of the data in the context of testing exchangeability, which we show may be viewed as the representative inversion for exchangeability.
1.6 Simulations, application and illustration
We empirically illustrate our methods in two simulation experiments and an application. The first simulation mimics a standard case-control experiment under random treatment allocation. In the second experiment, we compare a sign-flipping e-process to one based on de la Peña (1999), and find that ours is dramatically more powerful. The application is to the ‘hot hand’ phenomenon in basketball, which is the belief that hitting a basketball shot increases the chances of hitting subsequent shots (Gilovich et al., 1985). This is frequently studied by assuming that the shot outcomes are exchangeable in absence of the hot hand, so that rejecting exchangeability also rejects the hot hand (Miller and Sanjurjo, 2018). We leverage the powerful merging properties of e-values by multiplying e-values across players to obtain a more powerful e-value. This merging of evidence is highly relevant for the hot hand, as a single shot sequence of a player is known to contain little evidence regarding the hot hand (Ritzwoller and Romano, 2022).
In Appendix A, we illustrate these methods on the problem of testing invariance on $\mathbb{R}^{d}$, $d\geq 1$, under an arbitrary compact group of orthonormal matrices, against a simple alternative that is a location shift under normality. For the special case of spherical invariance, this is connected to an example from Lehmann and Stein (1949) regarding the optimality of the $t$-test, which we slightly generalize. We also consider sign-symmetry, which produces an e-value that can be viewed as an admissible version of an e-value based on de la Peña (1999). Furthermore, we consider exchangeability where we discover an interesting link to the softmax function.
1.7 Related literature
At first glance, our work may seem intimately related to the work of Pérez-Ortiz et al. (2024). However, they consider invariance of collections of distributions (both the null and the alternative), whereas we consider invariance of distributions themselves. Specifically, a collection of distributions $\mathcal{P}$ is said to be invariant under a transformation $G$ if for any $\mathbb{P}\in\mathcal{P}$, its transformation $G\mathbb{P}$ by $G$ is also in $\mathcal{P}$. In contrast, invariance of a distribution $\mathbb{P}$ means that its transformation $G\mathbb{P}$ is equal to $\mathbb{P}$ itself. Intuitively, their work can be interpreted as testing in the presence of an invariant model, whereas we consider testing the data generating process is invariant.
As our null hypothesis consists exclusively of invariant distributions it is technically also invariant, so that one may believe their results may still apply under appropriate assumptions on the alternative. However, Pérez-Ortiz et al. (2024) require that the group action is free, which means that if $G\mathbb{P}=\mathbb{P}$ for some $\mathbb{P}\in\mathcal{P}$ then $G$ must be the identity element. In other words, applying a non-identity transformation to $\mathbb{P}$ must change it. Our form of invariance instead requires that $G\mathbb{P}=\mathbb{P}$ for every $G$ and $\mathbb{P}\in\mathcal{P}$. This means that the settings do not overlap, except for the uninteresting setting that the group only contains the identity element.
Vovk (2025) independently derives a permutation e-value for testing exchangeability of binary random variables against a single specific alternative hypothesis. Moreover, Vovk (2021) explores testing exchangeability in a sequential setup, by passing from the original data to its ranks. He exploits the fact that the sequential ranks are independent from the past ranks under exchangeability. He then converts these ranks into independent e-values, which are multiplied together to construct a test martingale under the rank-filtration. Lardy and Pérez-Ortiz (2024) apply this rank-based approach to testing group invariance in a setting similar to Section 6 and Appendix B. We show how rank-based approaches may be viewed as a special case of using a unique representative inversion, where ranks appear as the representative inversion for exchangeability.
A link between the softmax function and e-values for exchangeability was also made in unpublished early manuscripts of Wang and Ramdas (2022) and Ignatiadis et al. (2023), which they call a ‘soft-rank’ e-value. In Remark 15 in Appendix A.6, we explore the connection to our softmax likelihood ratio statistic, and find that their soft-rank e-value can be interpreted as a more volatile version.
Testing the symmetry of a distribution, which we touch in Appendix A.7, was also studied by Ramdas et al. (2022a), Vovk and Wang (2024) and Larsson et al. (2024).
1.8 Notation and underlying assumptions
Throughout the paper, every ‘space’ we consider is assumed to be second-countable locally compact Hausdorff, equipped with a Borel $\sigma$-algebra. We intentionally suppress this topology and the $\sigma$-algebra whenever possible, for notational conciseness. To avoid ambiguity, we sometimes write expectations $\mathbb{E}$ with a superscript and/or subscript $\mathbb{E}_{X}^{\mathbb{P}}$ to make explicit the measure over which is being integrated ($\mathbb{P}$), and the random variables over which the integration takes place ($X$). We use similar subscripts for probabilities.
2 Background: group invariance
In this section, we discuss all the necessary background on group invariance. We recommend Eaton (1989) for deeper treatment of invariance in statistics.
2.1 Compact groups
A group $\mathcal{G}$ is a set equipped with some associative binary operator ‘$\times$’ that is closed under composition and inversion, and contains an identity element $I$. For brevity, we use juxtaposition $G_{1}G_{2}=G_{1}\times G_{2}$ to denote the binary operation, $G_{1},G_{2}\in\mathcal{G}$. A subset of a group that is also a group is called a subgroup.
Throughout, unless stated otherwise, all groups we consider are compact groups. Compact groups are groups that are also compact topological spaces. Compact groups are special in that they admit a unique invariant probability measure $\mathbb{P}_{\text{Haar}}$ on $\mathcal{G}$ called the Haar probability measure, which satisfies
$\mathbb{P}_{\text{Haar}}(GA)=\mathbb{P}_{\text{Haar}}(A),\text{ for all }G\in\mathcal{G},$
for every event $A$ on the group, where $GA:=\{Ga:a\in A\}$ is the event $A$ translated by $G$.
The Haar probability measure can be interpreted as the uniform probability measure on the group: whenever we shift an event $A$ by some element $G$ of the group, its probability remains unchanged. We use $\overline{G}$ to denote a Haar-distributed random variable on $\mathcal{G}$, and we also write $\textnormal{Unif}(\mathcal{G}):=\mathbb{P}_{\text{Haar}}$.
Example 1 (Orthonormal matrices).
A typical example of a compact group acting on a sample space is the collection of all $n\times n$ orthonormal matrices, $n\geq 1$, which acts on $\mathbb{R}^{n}$ through matrix multiplication. This group action rotates or flips $n$-vectors about the origin. Here, the identity element $I$ is the identity matrix. Moreover, the inverse of an orthonormal matrix is simply its transpose: $G^{-1}=G^{\prime}$, which is also orthonormal. The Haar measure is the uniform distribution over orthonormal matrices, and $\overline{G}$ is an orthonormal matrix drawn uniformly at random.
2.2 Group actions and orbits
In statistics, we are often interested in the action of a group $\mathcal{G}$ on a sample space $\mathcal{Y}$. We also denote such a group action through juxtaposition: $(G,y)\mapsto Gy$, and assume that it is continuous.
A group action partitions the sample space into orbits. The orbit of a sample point $y\in\mathcal{Y}$, denoted by $O_{y}=\{z\in\mathcal{Y}\ |\ z=Gy,\ G\in\mathcal{G}\}$, can be interpreted as the set of all sample points that can be reached when starting from $y$ and applying an element of the group to it. We use $\mathcal{Y}/\mathcal{G}$ to denote the collection of orbits. We assign a single point $[y]$ on each orbit as the orbit representative of the orbit $O_{y}$. That is, $[y]=Gy$ for some $G\in\mathcal{G}$. This means $O_{y}=O_{[y]}$ for any $y\in\mathcal{Y}$. We use $[\mathcal{Y}]$ to denote the collection of orbit representatives, and we call the map $[\cdot]:\mathcal{Y}\to[\mathcal{Y}]$ that maps $y$ to its orbit representative an orbit selector. We assume the orbit selector is chosen to be measurable, which is possible if $\mathcal{G}$ is compact.
Example 1 (Part B).
The group of orthonormal matrices acts on the sample space $\mathcal{Y}=\mathbb{R}^{n}$ through matrix multiplication. In this context, the collection of orbits $\mathcal{Y}/\mathcal{G}$ is the collection of hyperspheres in dimension $n$, each with a different radius. Given some unit vector $\iota$ we can assign the vector $r\iota$ as the orbit representative of the orbit with radius $r\geq 0$. The corresponding orbit selector is the map $Y\mapsto\|Y\|_{2}\iota$, since $Y$ is on the hypersphere with radius $\|Y\|_{2}$.
Example 2 (One orbit).
If there is just a single orbit, we say that the group acts transitively on the sample space. As the orbits partition the sample space, this means the sample space is the orbit.
This happens, for example, if our sample space is our group: $\mathcal{Y}=\mathcal{G}$. Another example is if we take Example 1 (Part B), but replace the sample space $\mathcal{Y}=\mathbb{R}^{n}$ with some hypersphere in dimension $n$. In Section 2.7, we discuss how we may generally reduce to a single orbit, which turns out to be a useful tool in statistical contexts.
2.3 Group invariant probability measures
In statistics, a sample space comes equipped with a collection of probability measures $\mathcal{P}$. The group action of $\mathcal{G}$ on $\mathcal{Y}$ induces a group action on the set of probabilities $\mathcal{P}$. We can define this group action $(G,\mathbb{P})\mapsto G\mathbb{P}$ as mapping the probability measure $\mathbb{P}$ to a probability measure $G\mathbb{P}$ that returns the $\mathbb{P}$-probability of the translation $G^{-1}A$ of an event $A$:
$G\mathbb{P}(A):=\mathbb{P}(G^{-1}A).$
We may then extend the idea of a Haar probability measure, which is an invariant probability on the group $\mathcal{G}$, to invariant probabilities on a sample space $\mathcal{Y}$: we say that a probability $\mathbb{P}$ is invariant if the group action does not affect the probability.
Definition 1 (Invariant probability).
$\mathbb{P}$ is an invariant probability measure if $G\mathbb{P}=\mathbb{P}$, for every $G\in\mathcal{G}$.
On each orbit $O$, there exists a unique invariant probability measure, which may therefore be safely called ‘the’ uniform probability $\textnormal{Unif}(O)$ on the orbit $O$. If there is just a single orbit, this means there is a single invariant probability. This happens, for example, if $\mathcal{Y}=\mathcal{G}$: the uniform probability is then the Haar measure $\textnormal{Unif}(\mathcal{G})$. If there are multiple orbits, there are generally multiple invariant probabilities: any probability mixture over uniform distributions on orbits is an invariant probability. In fact, in Lemma 1, we see that the converse also holds: any invariant probability may be viewed as a mixture over uniform probabilities on orbits.
Example 1 (Part C).
A typical example of an invariant distribution on $\mathbb{R}^{n}$ under the group of orthonormal matrices is the $n$-dimensional i.i.d. Gaussian distribution with mean zero and some variance $\sigma^{2}\geq 0$. In fact, this almost characterizes the Gaussian: the multivariate standard Gaussians are the only rotationally-invariant distributions that have i.i.d. marginals, by the Herschel-Maxwell theorem.
2.4 Equivalent characterizations of invariance
Beyond Definition 1, there exist several equivalent ways to characterize an invariant probability measure. In Lemma 1 we list a number of such equivalent definitions, expressed in terms of a random variable $Y$. These can also be expressed in terms of probability measures, if desired. But we find that discussing our results in the context of random variables generally yields more easily interpretable statements. A proof of these statements may be found in Chapter 4 of Eaton (1989).
Lemma 1 (Equivalent definitions of invariance).
$Y$ is an invariant random variable under a compact group $\mathcal{G}$ if one of the following equivalent conditions holds:
  1. 1.
    The law $\mathbb{P}_{Y}$ of $Y$ is invariant,
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
[…]
  5. 5.
    The conditional law of $Y$ given $O_{Y}$ is $\textnormal{Unif}(O_{Y})$,
where condition 5 may be read as ‘there exists a version of the conditional law’, to acknowledge the fact that such a conditional law may only be defined up to $\mathbb{P}_{[Y]}$-almost sure equivalence.
Condition 4 is particularly insightful: it states that a draw from an invariant random variable $Y$ can be decomposed (deconvolved) into first sampling an orbit representative $[Y]$ and subsequently multiplying it by $\overline{G}$, independently sampled uniformly from $\mathcal{G}$.
Condition 5 restates this in terms of orbits: a draw from $Y$ can be viewed as first sampling an orbit $O_{Y}$ using some unspecified process, and subsequently sampling uniformly from this orbit. This decomposition is the key to testing invariance, where the idea is to effectively discard the first part of this sampling process, and only test whether $Y$ is uniform conditional on the orbit in which it is observed.
[…]
Example 1 (Part C).
A draw from an $n$-dimensional standard Gaussian $Y$ can be decomposed into first sampling a radius (and so orbit) from a $\chi_{n}$-distribution, and subsequently independently drawing an $n$-vector uniformly from the hypersphere with this radius (from the sampled orbit). Lemma 1 states that any rotationally invariant random variable can be characterized as such: first sampling a radius using some distribution, and subsequently independently drawing an $n$-vector uniformly from the sampled orbit.
2.5 Constructing invariant probability measures and random variables
It is possible to create an invariant probability measure out of any probability measure $\mathbb{P}$, by “averaging” it over the group:
¯$\overline{\mathbb{P}}:=\mathbb{E}_{\overline{G}}[\overline{G}\mathbb{P}],$
where $\overline{G}\sim\textnormal{Unif}(\mathcal{G})$. These group-averaged invariant measures play a central role in optimal tests and e-values. For an invariant probability $\mathbb{P}$, we have $\mathbb{P}=\overline{\mathbb{P}}$, so that this averaging has no effect. We can also express this in terms of random variables: if $Y\sim\mathbb{P}$, then $\overline{G}Y\sim\overline{\mathbb{P}}$.
2.6 Invariance through a statistic
Sometimes, we only look at our random variable $Y$ through a statistic $S$, such as a test statistic. In such situations, it does not matter whether $Y$ is actually invariant; it only matters whether it looks invariant when viewed through this statistic. This leads us to the following weaker notion of invariance, which recovers the standard notion if $S$ is invertible. This will yield a more general condition under which tests for group invariance are valid, given the choice of test statistic.
Definition 2 (Invariance through a statistic).
[…]
We illustrate the difference between invariance and invariance through a statistic in two examples. In Example 3, the random variable is invariant. In Example 4, the underlying random variable is not invariant, but looks invariant through certain statistics.
Example 3 (Invariance under permutations: exchangeability).
Suppose we have a single bag and fill it with the numbers 1, 2, 3 and 4. We now sample uniformly without replacement from this bag and arrange the numbers in the order they were drawn. As each order has the same probability, we say that this outcome is exchangeable: invariant under all permutations of the numbers $\{1,2,3,4\}$.
Example 4 (Not exchangeable, but exchangeable through a statistic).
Suppose we have two bags. We fill one with the numbers 1 and 2, and the other with numbers 3 and 4. We start by picking a bag with equal probability, and then sequentially draw both numbers from the bag in an exchangeable manner. Next, we take the other bag and do the same, after which we arrange the numbers in the order they were drawn.
Here, the resulting order of the numbers is not invariant under all permutations: out of 24 permutations, only the 8 orders 1234, 1243, 2134, 2143, 3412, 3421, 4312 and 4321 can occur. The order does look exchangeable through the statistic that returns only the first position. Indeed, every number is equally likely to land in the first position both under our sampling process, and if we had used an exchangeable sampling process.
The order also looks exchangeable through the statistic $S$ that returns the relative ranks of the first two positions: $S(12\cdot\cdot)=12$, $S(21\cdot\cdot)=21$, $S(34\cdot\cdot)=12$, $S(43\cdot\cdot)=21$. This is because the ranks 12 and 21 happen with equal probability both under our sampling process, and under an exchangeable sampling process.
While we only consider a single orbit in this example — namely the permutations of 1234 — the example extends to multiple orbits. Indeed, we may view the numbers 1234 as determined by some preceding sampling process.
Remark 1.
[…]
2.7 Reducing to a single orbit and representative inversion kernels
When testing invariance under a group of permutations (exchangeability), it is common to convert data to its ranks (relative to a canonical ordering). Similarly, it is common to only look at the signs of random variables that are invariant under sign-flipping (symmetric about zero) and a normalized statistic $y\mapsto y/\|y\|_{2}$ for rotation-invariant random variables. Functions of these statistics give rise to rank tests, sign tests and the $t$-test (see Example 5).
We believe the underlying reason for the popularity of these statistics is the fact that the null distribution of these statistics is the same on each orbit. As a consequence, we may pick an arbitrary orbit and behave as if we are testing whether the random variable is uniform on this arbitrary orbit, which is generally a much simpler task.
We find that passing to ranks, signs and normalization can be viewed as a special case of a more general recipe, as detailed in Remark 2. In particular, we may construct a statistic with the desired property by means of a (representative) inversion kernel $\gamma$, which maps the data to the group $\mathcal{G}$ (see Kallenberg (2011) and Chapter 7 of Kallenberg (2017)). The key property of an inversion kernel for testing group invariance is that if $Y$ is $\mathcal{G}$ invariant, then $\gamma(Y)\sim\textnormal{Unif}(\mathcal{G})$, independently of the orbit in which $Y$ lands. Inversion kernels were first used in the context of group invariance testing by Chiu and Bloem-Reddy (2023).
To define an inversion kernel, it is convenient to first assume that the group $\mathcal{G}$ acts freely on $\mathcal{Y}$. This means that $Gy=y$ for some $y\in\mathcal{Y}$ implies $G=I$. In the context of permutations, this assumption means that there are no ties: indeed, barring ties, a non-identity permutation of data always modifies the original data. Under this assumption, we can uniquely define the inversion kernel as a map $\gamma:\mathcal{Y}\to\mathcal{G}$ that takes $y$ and returns the element $G$ that carries the representative element $[y]$ on the orbit of $y$ to $y$. That is, $\gamma(y)[y]=y$.
If the group action is not free, then there may exist multiple elements in $\mathcal{G}$ that carry $[y]$ to $y$, so that $\gamma(y)$ is not uniquely defined. For the non-free setting, we overload the notation of $\gamma$ so that $\gamma(y)$ is uniformly drawn from the elements in $\mathcal{G}$ that carry $[y]$ to $y$, which is well-defined as shown in Theorem 7.14 of Kallenberg (2017). This gives us $\gamma(y)[y]=y$ almost surely. Section D in the Supplementary Material contains a concrete illustration of a setting where $\gamma$ is randomized in this manner, and an intuition of why it is possible to construct a uniform draw from such elements.
Remark 2 (Relationship inversion kernel and ranks, signs and normalization).
Ignoring ties, the relationship between the inversion kernel and ranks is that ranks are in bijective correspondence to the group of permutations. In case of ties, the inversion kernel can be viewed as a slight generalization, that breaks ties through randomization by smearing out the probability mass over different permutations that yield the same data due to ties. Such randomized tie-breaking is already used in many applications of rank-based testing.
Similarly, barring zeros, the signs of a tuple of data $(X_{1},...,X_{n})$ are in bijective correspondence to a group of sign-flips $\{-1,1\}^{n}$. In $\mathbb{R}^{2}\setminus\{0\}$, the normalized vector $y\mapsto y/\|y\|_{2}$ is in bijective correspondence to the special orthogonal group of ‘rotations’. In higher dimensions, there may be multiple rotations that carry the representative element $[y/\|y\|_{2}]$ to $y/\|y\|_{2}$, and the resulting inversion kernel can be interpreted as uniformly sampling one of these rotations.
3 Tests and e-values for group invariance
3.1 Hypothesis and e-value
Our goal is to measure evidence against the hypothesis that a random variable $Y$ is drawn from some $\mathcal{G}$ invariant distribution:
$Y\text{ is }\mathcal{G}\text{ invariant}.$
Equivalently, we test whether the latent distribution from which $Y$ is sampled is in the collection $H:=\{\mathbb{P}:\mathbb{P}\textnormal{ is }\mathcal{G}\textnormal{ invariant}\}$. For this purpose, we use an e-value $\varepsilon:\mathcal{Y}\to[0,\infty]$, which is said to be valid for the hypothesis $H$ if
$\sup_{\mathbb{P}\in H}\mathbb{E}^{\mathbb{P}}\,\varepsilon\leq 1.$
We say an e-value is exact for $H$ if $\mathbb{E}^{\mathbb{P}}\,\varepsilon=1$ for every $\mathbb{P}\in H$.
Remark 3 (Exact e-value).
The term ‘exact e-value’ with respect to a hypothesis $H$ is typically reserved for the property $\sup_{\mathbb{P}\in H}\mathbb{E}^{\mathbb{P}}\,\varepsilon=1$, which is weaker than what we call exact here: $\mathbb{E}^{\mathbb{P}}\,\varepsilon=1$ for every $\mathbb{P}\in H$. Our property may be viewed as ‘uniformly exact’ over the hypothesis consisting of all invariant probabilities on our sample space, but for brevity we simply refer to it as exact.
3.2 Characterizing e-values for group invariance
We immediately present our first result, which characterizes valid and exact e-values for group invariance. It states that an e-value is valid for group invariance if and only if it is valid for a uniform distribution on each orbit. A formal proof is provided in Appendix E.1.
Theorem 1.
Let $\varepsilon:\mathcal{Y}\to[0,\infty]$. Then,
  • (i)
    $\varepsilon$ is a valid e-value for $\mathcal{G}$ invariance if and only if $\mathbb{E}^{\textnormal{Unif}(O)}\,\varepsilon\leq 1$, for every $O\in\mathcal{Y}/\mathcal{G}$,
  • (ii)
    $\varepsilon$ is an exact e-value for $\mathcal{G}$ invariance if and only if $\mathbb{E}^{\textnormal{Unif}(O)}\,\varepsilon=1$, for every $O\in\mathcal{Y}/\mathcal{G}$.
Note that $\overline{G}y\sim\textnormal{Unif}(O)$ for every $y\in O$. Hence, validity may be equivalently formulated as $\mathbb{E}_{\overline{G}}[\varepsilon(\overline{G}y)]\leq 1$ for every $y\in\mathcal{Y}$.
As the orbits partition the sample space, our data $Y$ lands in exactly one orbit: $O_{Y}$. As a consequence, we actually only need our e-value to be valid for $\textnormal{Unif}(O_{Y})$, conditional on $O_{Y}$, as captured in Corollary 1. This is the key that facilitates testing group invariance. In particular, we may view the problem of testing group invariance as first observing the orbit $O_{Y}$ and then testing the simple hypothesis that $Y$ is uniform on $\textnormal{Unif}(O_{Y})$.
Corollary 1.
$\varepsilon:\mathcal{Y}\to[0,\infty]$ is a valid (exact) e-value for $\mathcal{G}$ invariance if and only if it is a valid (exact) e-value for $\textnormal{Unif}(O_{Y})$, conditionally on $O_{Y}$.
Remark 4 (Non-compact groups).
For a possibly non-compact group $\mathcal{G}$ acting on a Borel space, we may still obtain a characterization as in condition 5 in Lemma 1. Indeed, a $\mathcal{G}$ invariant probability may be viewed as a mixture over ergodic probabilities (see Kallenberg (2021) Theorem 25.24). We may use this fact to obtain a result analogous to Theorem 1, with the ergodic probabilities taking the place of the uniform distributions on the orbits $\textnormal{Unif}(O)$. Unfortunately, Corollary 1 falls apart, as these ergodic probabilities may have overlapping support: they are supported on unions of orbits. This means that by observing the data $Y$, we generally cannot identify a single ergodic probability in the way that we can identify $\textnormal{Unif}(O_{Y})$ in the compact setting. This property is key for many of our results, so that we stick to compact groups throughout, and only briefly return to non-compact groups in the sequential setting in Remark 11.
Remark 5 (Infimum representation).
An alternative way to represent a valid e-value for group invariance is as an infimum over orbit-wise e-values. Indeed, we may construct an e-value $\varepsilon^{O}:\mathcal{Y}\to[0,\infty]$ for each orbit, satisfying
$\mathbb{E}^{\textnormal{Unif}(O)}\,\varepsilon^{O}\leq 1,$ (2)
and $\varepsilon^{O}(y)=\infty$ if $y\not\in O$. We can then view the e-value for $\mathcal{G}$ invariance as $\varepsilon(y)=\textnormal{ess\,inf}_{O}\,\varepsilon^{O}(y)$. While this may seem superfluous, this representation is useful in the sequential setting in Section 5, where we cannot simply condition on the orbit, as the orbit may not be measurable at the moment of testing. This is effectively the same problem we encounter with ergodic probabilities in Remark 4. There, e-values for non-compact groups may also be represented as an infimum over ergodic-wise e-values analogous to (2).
3.3 Generic traditional tests for group invariance
Traditionally, evidence against group invariance is measured through a ‘group invariance test’. Such a test may be viewed as a special case of an e-value that either emits zero evidence or $1/\alpha$ evidence against the hypothesis, for some pre-specified significance level $\alpha\in(0,1)$. We cover this special e-value here first and move to more general e-values in the next section.
Given any test statistic $T:\mathcal{Y}\to\mathbb{R}$, ideally designed to be large under the alternative, the traditional test for group invariance $\varepsilon_{\alpha}:\mathcal{Y}\to[0,1/\alpha]$ is given by
$\varepsilon_{\alpha}(y)=\tfrac{1}{\alpha}\mathbb{I}\left\{T(y)>q_{\alpha}^{\overline{G}}[T(\overline{G}y)]\right\}+\tfrac{c([y])}{\alpha}\mathbb{I}\left\{T(y)=q_{\alpha}^{\overline{G}}\left[T(\overline{G}y)\right]\right\},$ (3)
[…]
It is easy to show and well-known that $\mathbb{E}_{\overline{G}}\,\varepsilon_{\alpha}(\overline{G}y)=1$ by plugging in the appropriate choice for $c([y])$. By Theorem 1, this means the test is exact regardless of the choice of test statistic $T$. This property is a well-known feature of group invariance tests, and the main driver behind their popularity.
We slightly extend this by showing it also holds if $Y$ is $\mathcal{G}$ invariant through $T$, as in Definition 2. To the best of our knowledge, this is novel. Moreover, the result actually holds conditional on the orbit of $Y$. As a consequence, we may for example choose the test statistic based on the orbit, if desired. Its proof can be found in Section E.2 in the Supplementary Material.
Theorem 2.
If $Y$ looks $\mathcal{G}$ invariant through $T$, then $\mathbb{E}_{\overline{G}}\,\varepsilon_{\alpha}(\overline{G}Y)=1$.
The $t$-test, which is an example of a group invariance test, is given in Example 5. Example 6 covers the most basic form of conformal inference (Shafer and Vovk, 2008; Angelopoulos et al., 2025).
Example 5 ($t$-test).
Suppose $\mathcal{Y}=\mathbb{R}^{n}$ and $T$ is defined as $T(y)=\iota^{\prime}y/\|y\|_{2}$, where $\iota$ is some unit vector, typically $\iota=n^{-1/2}(1,1,\dots,1)$. If $Y$ is spherically invariant through $T$, then $T(Y)$ is Beta$(\tfrac{n-1}{2},\tfrac{n-1}{2})$-distributed on $[-1,1]$ (see e.g. Koning and Hemerik (2023) for a proof) conditional on every orbit, and so unconditionally as well. Equivalently, $\sqrt{n-1}T(Y)/\sqrt{1-T(Y)^{2}}$ is $t$-distributed. The resulting test for spherical invariance is also known as the $t$-test.
Example 6 (Conformal inference).
Suppose $\mathcal{Y}=\mathbb{R}^{n+1}$ and $\mathcal{G}$ is the group of permutations acting on the canonical basis of $\mathbb{R}^{n+1}$. Let $Y^{n+1}$ be a $\mathcal{G}$ invariant (exchangeable) random variable on $\mathcal{Y}$, and let $T:\mathcal{Y}\to\mathbb{R}$ be a test statistic that only depends on the final element $Y_{n+1}$. Suppose we only observe $Y^{n}=(Y_{1},\dots,Y_{n})$ and want to test whether the unobserved $Y_{n+1}$ equals $y^{*}$. We can then use the permutation test based on $T((Y^{n},y^{*}))$, which is also known as conformal inference. Repeating this test for all $y^{*}\in\mathcal{Y}$ and collecting the values of $y^{*}$ for which we do not reject yields the conformal prediction set, which is a confidence set for $Y_{n+1}$ in $\mathbb{R}$.
3.4 Generic e-values for group invariance
We now show how we may move beyond traditional group invariance tests, by deriving ‘generic’ exact e-values for group invariance based on some test statistic $T$.
As with the traditional group invariance tests in Section 3.3, we retain great freedom in our selection of the test statistic. In particular, let $T$ be some arbitrary non-negative test statistic that is appropriately integrable on every orbit $O\in\mathcal{Y}/\mathcal{G}$. Specifically, we require $0<\mathbb{E}_{\overline{G}}T(\overline{G}y)<\infty$ for every $y\in\mathcal{Y}$.
Based on this test statistic $T$, we consider the e-value
$\varepsilon_{T}(y)=\frac{T(y)}{\mathbb{E}_{\overline{G}}T(\overline{G}y)},$ (4)
where $\overline{G}\sim\textnormal{Unif}(\mathcal{G})$. The interpretation is that $\varepsilon_{T}(Y)$ is large if $T(Y)$ is large compared to its average value on the orbit of $Y$.
Theorem 3 shows that this e-value is exact, and that any exact e-value for $\mathcal{G}$ invariance may be construed in this manner. The proof leverages Theorem 1, and may be found in Appendix E.3.
Theorem 3.
An e-value $\varepsilon$ is exact for $\mathcal{G}$ invariance if and only if it is of the form $\varepsilon_{T}$ for some statistic $T$.
By Theorem 3, we can use any appropriately integrable test statistic $T$ to construct an exact e-value for $\mathcal{G}$ invariance. In fact, as a non-exact e-value is such a statistic, we can plug it in for $T$ to transform it into an exact e-value. We exploit this trick in Appendix A.7.
Proposition 1 shows that we may also relax the assumption of $\mathcal{G}$ invariance by incorporating the choice of the test statistic. This assumption is weaker than $\mathcal{G}$ invariance through $T$, which we assume for Proposition 2: we only require the expectation of $T(Y)$ on each orbit to equal the uniform orbit-average $\mathbb{E}_{\overline{G}}T(\overline{G}[y])$. Its proof is given in Appendix E.4.
Proposition 1.
Assume that $\mathbb{E}_{Y}[T(Y)]=\mathbb{E}_{\overline{G}}[T(\overline{G}[Y])]$ conditional on $O_{Y}$, where $\overline{G}\sim\textnormal{Unif}(\mathcal{G})$, independently. Then $\varepsilon_{T}$ is a valid e-value for $\mathcal{G}$-invariance.
Example 7 (An e-value version of the $t$-test).
Continuing from Example 5, one may desire to derive “the e-value version” of the $t$-test. But because e-values are a (rich) generalization of binary tests, there is no unique e-value version of the $t$-test: any e-value $\varepsilon_{T}$ based on a statistic $T$ that is non-decreasing in $\iota^{\prime}y/\|y\|_{2}$ could reasonably qualify.
Example 8 (Conformal inference with e-values).
Continuing the setup from Example 6, if $T$ is a non-negative test statistic that only depends on the final element, then $T((Y^{n},y^{*}))/\mathbb{E}_{\overline{G}}T(\overline{G}(Y^{n},y^{*}))$ is an exact e-value for conformal inference.
3.5 Obtaining the normalization constant and Monte Carlo group invariance e-values
The main computational challenge when using e-values for group invariance is the computation of the normalization constant $\mathbb{E}_{\overline{G}}T(\overline{G}Y)$. As the group $\mathcal{G}$ is often large, simply averaging $T(GY)$ over all $G$ may not be feasible. However, the normalization constant can be estimated.
We borrow some ideas from traditional group invariance testing. The simplest idea is to use a Monte Carlo approach by replacing $\overline{G}$ with a random variable $\overline{G}^{M}$ that is uniformly distributed on a collection $(\overline{G}^{(1)},\overline{G}^{(2)},\dots,\overline{G}^{(M)})$ of $M\geq 1$ mutually independent and identically distributed copies of $\overline{G}$, independent from $Y$. Writing $\overline{G}^{(0)}=I$, this yields the Monte Carlo group invariance e-value
$\varepsilon_{T}^{M}(y)=\frac{T(y)}{\frac{1}{M+1}\sum_{i=0}^{M}T(\overline{G}^{(i)}y)}.$
This Monte Carlo group invariance e-value is exact in expectation over the Monte Carlo sample, as captured in Theorem 4. The proof can be found in Appendix E.5, and relies on establishing the exchangeability of $T(\overline{G}^{(0)}Y),\dots,T(\overline{G}^{(M)}Y)$ under the null hypothesis, and then applying Theorem 3.222We thank an anonymous referee for suggesting this proof strategy.
Theorem 4.
The Monte Carlo e-value $\varepsilon_{T}^{M}$ is exact in expectation over the Monte Carlo samples: $\mathbb{E}_{\overline{G}^{(1)},\dots,\overline{G}^{(M)}}\mathbb{E}_{\overline{G}}[\varepsilon_{T}^{M}(\overline{G}y)]=1$, for every $y\in\mathcal{Y}$.
While the resulting e-value is exact in expectation for any number of Monte Carlo draws $M$, a larger number of draws should generally improve the estimation of the normalization constant $\mathbb{E}_{\overline{G}}T(\overline{G}Y)$, and thereby reduce the resampling risk of the Monte Carlo e-value: the sensitivity of $\varepsilon_{T}^{M}$ to the drawn sample.
Recently, Fischer and Ramdas (2025) studied sequential Monte Carlo testing in the context of traditional Monte Carlo group invariance tests, where the number of Monte Carlo draws $M$ need not be pre-specified but may be determined adaptively as a stopping time. While determining $M$ adaptively should generalize to Monte Carlo e-values, the approach of Fischer and Ramdas (2025) is tailored towards making a binary decision, and we believe it would require substantial modification to apply here. This is highlighted by Stoepker and Castro (2024) in the context of p-values, who argue that a much larger number of Monte Carlo draws is typically desirable for a continuous measure of evidence when compared to a binary decision. We believe there is much to explore in the context of sequentially drawn Monte Carlo e-values, but leave this topic for future work.
Instead of Monte Carlo sampling, a different approach that has been explored in traditional group invariance testing is to replace $\overline{G}$ in (4) with a random variable that is uniformly distributed on a compact subgroup of $\mathcal{G}$ (Chung and Fraser, 1958). As invariance under $\mathcal{G}$ implies invariance under every subgroup, this still guarantees the resulting e-value is valid. Such a subgroup may also be easier to work with than $\mathcal{G}$ itself, if applied inside multiple testing procedures. Moreover, Koning and Hemerik (2023) note that we can actually strategically select the subgroup based on the test statistic and alternative, and select a subgroup that yields high power. Koning (2024b) observes that this can even yield testing methods that are more powerful than if we use the entire group $\mathcal{G}$. This approach can be combined with Monte Carlo sampling by sampling from the subgroup. Ideas to go beyond uniform distributions on subgroups appear in Hemerik and Goeman (2018) and Ramdas et al. (2023a).
If $T(Y)$ has the same distribution on each orbit $O\in\mathcal{Y}/\mathcal{G}$ under the null hypothesis, then the distribution can even be easily precomputed, as it is not necessary to know the orbit of the data $Y$. This is true for statistics that are a function of an inversion kernel, as in the case of sign- and rank-tests.
4 Optimal e-values for group invariance
While Section 3 describes a flexible way to construct an e-value for group invariance, it does not instruct us how to choose a good e-value for testing group invariance. In this section we provide guidance, by deriving optimal e-values for group invariance.
4.1 Background: optimality objectives for e-values
To speak about an ‘optimal’ e-value, we must establish an objective under which it is optimal. To do this, it is typical to select some alternative distribution $\mathbb{Q}$ under which our e-value should be ‘large’ in some expected sense. In this section, we follow the unification between optimal traditional testing and optimal e-values developed in Koning (2024c).
We can view traditional Neyman–Pearson-style testing as using an e-value that maximizes the ‘power’ $\mathbb{E}^{\mathbb{Q}}[\varepsilon\wedge 1/\alpha]$. For a simple null $\mathbb{P}$, the Neyman–Pearson lemma tells us this target is optimized by
$\varepsilon^{*}=\begin{cases}1/\alpha,&\text{ if }\frac{q}{p}>c_{\alpha},\\ k,&\text{ if }\frac{q}{p}=c_{\alpha},\\ 0&\text{ if }\frac{q}{p}<c_{\alpha},\end{cases}$ (5)
where $p$ and $q$ are densities of $\mathbb{P}$ and $\mathbb{Q}$ with respect to some reference measure, which always exists (e.g. $(\mathbb{P}+\mathbb{Q})/2$), and $k\in[0,1/\alpha]$ and $c_{\alpha}\geq 0$ are some constants. Here, a value $\varepsilon=1/\alpha$ corresponds to the traditional notion of a rejection at level $\alpha$, $\varepsilon=0$ as a non-rejection, and $\alpha\varepsilon$ is traditionally viewed as a probability with which one should subsequently reject at level $\alpha$ using external randomization.
In Koning (2024c), it is argued that e-values are merely a fuzzy or ‘continuous’ interpretation of a test. At the same time, Neyman–Pearson-style optimal e-values hardly use the interval $[0,1/\alpha]$: the middle case when $q/p=c_{\alpha}$ happens with zero or small probability in many settings. Hence, we must move to different power objectives to truly move away from the traditional binary form of testing.
The most popular objective in the e-value literature, sometimes even called ‘e-power’, is the expected logarithm under the alternative (Shafer, 2021; Koolen and Grünwald, 2022; Grünwald et al., 2024): $\mathbb{E}^{\mathbb{Q}}[\log(\varepsilon)]$. We call e-values that optimize this target ‘log-optimal’. Log-optimal e-values have recently been shown to exist regardless of the hypothesis, and so must also exist when testing group invariance (Larsson et al., 2024).
[…]
$\varepsilon^{*}(y)=(U^{\prime})^{-1}\left(\lambda^{*}\frac{p(y)}{q(y)}\right),$ (6)
for some appropriate normalization constant $\lambda^{*}$.
While expected utility certainly does not exhaust all possible optimization objectives, this captures a rich palette of objectives to choose from.
4.2 Characterizing e-values and optimal e-values for group invariance
While e-values for simple null hypotheses as in (6) are well-understood, optimal e-values for composite hypotheses are significantly more challenging to characterize. In Theorem 5, we present the key result for deriving optimal e-values for $\mathcal{G}$ invariance. Its proof is presented in Appendix E.6.
The result relies on splitting the optimization problem into ‘local’ optimization problems on each orbit. In particular, we start by deriving a ‘locally’ optimal e-value $\varepsilon_{|O}^{*}$ for $\textnormal{Unif}(O)$ on each orbit $O\in\mathcal{Y}/\mathcal{G}$. We then stitch these locally optimal e-values together by setting the optimal e-value $\varepsilon^{*}(y)$ equal to the value of $\varepsilon_{|O_{[y]}}^{*}(y)$ on the orbit $O_{[y]}$ in which $y$ is observed: $\varepsilon^{*}(y)=\varepsilon_{|O_{[y]}}^{*}(y)$. This is useful, because finding an optimal e-value $\varepsilon_{|O}^{*}$ on some orbit $O$ is a much simpler task: this comes down to testing the simple hypothesis $\textnormal{Unif}(O)$ on the orbit.
To allow for such an orbit-by-orbit strategy, we must assume that we are maximizing an objective $K$ that may be decomposed into local objectives $K_{O}$ on each orbit $O$. This condition seems quite mild: it holds, for example, for the expected utility-type objectives discussed in Section 4.1.
To present such an objective, we need to introduce some concepts. In order to present the result, let $F_{+}^{\mathcal{A}}$ denote the set of $[0,\infty]$-valued measurable functions on the space $\mathcal{A}$. For a measurable subspace $\mathcal{B}\subseteq\mathcal{A}$, we use $f_{|\mathcal{B}}$ to denote the restriction of $f\in F_{+}^{\mathcal{A}}$ to $\mathcal{B}$. Note that $f_{|\mathcal{B}}\in F_{+}^{\mathcal{B}}$. We use $K:F_{+}^{\mathcal{Y}}\to[0,\infty]$ to denote the aggregate objective, which we define by
$K(f)=\Psi\left(\bigl{(}K_{O}\bigl{(}f_{|O}\bigr{)}\bigr{)}_{O\in\mathcal{Y}/\mathcal{G}}\right),$ (7)
where $K_{O}:F_{+}^{O}\to[0,\infty]$ is an orbit-based objective, for each orbit $O\in\mathcal{Y}/\mathcal{G}$, and the aggregating function $\Psi:[0,\infty]^{\mathcal{Y}/\mathcal{G}}\to[0,\infty]$ is non-decreasing in each of its inputs.
Theorem 5 (Local optimality $\implies$ global optimality).
Let $\varepsilon^{*}\in F_{+}^{\mathcal{Y}}$. Suppose that for each $O\in\mathcal{Y}/\mathcal{G}$:
  • (i)
    $\varepsilon_{|O}^{*}$ is a valid e-value for $\mathrm{Unif}(O)$,
  • (ii)
    $K_{O}\left(\varepsilon_{|O}^{*}\right)\geq K_{O}(\varepsilon)$ for every e-value $\varepsilon\in F_{+}^{O}$ that is valid for $\mathrm{Unif}(O)$.
Then $\varepsilon^{*}$ is valid for $\mathcal{G}$ invariance and $K$-optimal: $K(\varepsilon^{*})\geq K(\varepsilon)$ for every e-value $\varepsilon$ that is valid for $\mathcal{G}$ invariance.
Remark 6 (Optimal uniformly in aggregation functions).
An implication of Theorem 5 is that an e-value that is optimal for every $K_{O}$, $O\in\mathcal{Y}/\mathcal{G}$, is also optimal for any choice of aggregation function $\Psi$. We exploit this idea in Section 4.4.
Remark 7 (Beyond group invariance).
The result also goes through if the partitioning of the sample space is not generated by a group, and the null hypothesis is simply some probability mixture over known distributions on each of the subsets in the partition. Our group structure naturally generates such a partition, where the distribution on each subset in the partition is uniform. This approach may be interesting to derive optimal e-values in other contexts, but we did not explore this further.
4.3 Expected utility-optimal e-values for group invariance: alternative on sample space
We now apply Theorem 5 to derive optimal e-values for group invariance for several expected-utility-type power objectives. A remarkable feature of these optimal e-values is that they may be expressed in terms of unconditional densities, instead of conditional densities on orbits. This is a feature of the fact that such optimal e-values are (non-decreasing) functions of likelihood ratios (Koning, 2024c), combined with the fact that they share the same mixing distribution which drops out in the likelihood ratio.
For an alternative $\mathbb{Q}$ on $\mathcal{Y}$, recall from Section 2.5 that we use $\overline{\mathbb{Q}}$ to denote its group-averaged form. Let $q$ denote a density of $\mathbb{Q}$ with respect to some reference measure $\mathbb{H}$ that dominates both $\mathbb{Q}$ and $\overline{\mathbb{Q}}$. Such a reference measure exists: (¯$(\overline{\mathbb{Q}}+\mathbb{Q})/2$, but sometimes other choices of reference measure are more convenient; see Proposition 2.
We start with Theorem 6, which shows how our strategy captures and slightly refines the main result of Lehmann and Stein (1949) who derive Neyman–Pearson-style optimal e-values (tests) which correspond to the objective $\mathbb{E}^{\mathbb{Q}}[\varepsilon\wedge 1/\alpha]$. [^4] They seem to implicitly assume the existence of a $\mathcal{G}$ invariant reference measure: see Proposition 2. Its proof is presented in Appendix E.7. We use the convention $x/0=\infty$ for $x>0$, and $x/\infty=0$ for $x\geq 0$.
Theorem 6 (Neyman–Pearson-optimal e-values).
Fix $\alpha\in(0,1)$. A Neyman–Pearson-optimal e-value for $\mathcal{G}$ invariance against $\mathbb{Q}$ is given by
$\varepsilon^{\textnormal{NP}}(y)=\begin{cases}1,&\text{ if }q(y)=\overline{q}(y)=0,\\ 1/\alpha,&\text{ if }q(y)/\overline{q}(y)>c_{\alpha}^{[y]},\\ k_{[y]},&\text{ if }q(y)/\overline{q}(y)=c_{\alpha}^{[y]},\\ 0,&\text{ if }q(y)/\overline{q}(y)<c_{\alpha}^{[y]},\end{cases}$
for some orbit-dependent constant $k_{[y]}\in[0,1/\alpha]$, and critical value $c_{\alpha}^{[y]}\geq 0$ that ensure the e-value is exact on each orbit.
In Theorem 6, we have some flexibility in choosing $\varepsilon^{\textnormal{NP}}$ on null sets. For example, we may choose $\varepsilon^{\textnormal{NP}}(y)=\infty$ if $\overline{q}(y)=0$ and $q(y)>0$.
In Theorem 7 we present the expected utility-optimal e-value, for the target $\mathbb{E}^{\mathbb{Q}}[U(\varepsilon)]$, where $U$ is assumed to satisfy the regularity conditions described in Section 4.1. We omit its proof, as it is effectively the same as that of Theorem 6 but replacing the Neyman–Pearson lemma with the expected utility-optimal e-value result in Koning (2024c).
Theorem 7 (Utility-optimal e-values).
A $U$-optimal e-value for $\mathcal{G}$ invariance against $\mathbb{Q}$ is
$\varepsilon^{U}(y)=(U^{\prime})^{-1}\left(\lambda_{[y]}^{*}\frac{\overline{q}(y)}{q(y)}\right),$
if $q(y)>0$ or $\overline{q}(y)>0$, for some orbit-dependent normalization constant $\lambda_{[y]}^{*}$. Moreover, $\varepsilon^{U}(y)=1$ if $q(y)=\overline{q}(y)=0$.
Given the popularity of log-optimal e-values, we present this case separately as a corollary. Here, the normalization constant can be easily given explicitly.
Corollary 2 (Log-optimal).
A log-optimal e-value is given by
$\varepsilon^{\textnormal{log}}(y)=\left.\frac{q(y)}{\overline{q}(y)}\middle/\mathbb{E}_{\overline{G}}\left[\frac{q(\overline{G}y)}{\overline{q}(\overline{G}y)}\right]\right.,$
if $q(y)>0$ or $\overline{q}(y)>0$, and $\varepsilon^{\textnormal{log}}(y)=1$, otherwise.
The log-optimal e-value inherits its clean form from the generalized-means-optimal e-value, which maximizes $(\mathbb{E}^{\mathbb{Q}}[\varepsilon^{h}])^{1/h}$, $h\leq 1$, $h\neq 0$, and $\exp\left\{\mathbb{E}^{\mathbb{Q}}[\log(\varepsilon)]\right\}$ for $h=0$. This generalized-mean optimal e-value provides a simple parameter $h$ to tune the ‘riskiness’ of the e-value, where $h$ closer to 1 yields a more all-or-nothing-style e-value, whereas $h\to-\infty$ yields the constant e-value.
Corollary 3 (Generalized-means-optimal).
Suppose $\mathbb{E}_{\overline{G}}\left[\left(\frac{q(\overline{G}y)}{\overline{q}(\overline{G}y)}\right)^{\frac{1}{1-h}}\right]<\infty$. Then, an $h$-generalized-mean optimal e-value is given by
$\varepsilon^{(h)}(y)=\left.\left(\frac{q(y)}{\overline{q}(y)}\right)^{\frac{1}{1-h}}\middle/\mathbb{E}_{\overline{G}}\left[\left(\frac{q(\overline{G}y)}{\overline{q}(\overline{G}y)}\right)^{\frac{1}{1-h}}\right]\right.,$
if $q(y)>0$ or $\overline{q}(y)>0$, and $\varepsilon^{(h)}(y)=1$, otherwise.
Remark 8 (Link to ‘generic’ e-values).
These corollaries give guidance to the choice of test statistic $T$ for the ‘generic’ e-value presented in Section 3.4. Indeed, $T\propto q/\overline{q}$ and $T\propto(q/\overline{q})^{\frac{1}{1-h}}$ yield log-optimal and generalized-mean optimal e-values.
In case a $\mathcal{G}$ invariant reference measure is available, it is convenient to choose $q$ with respect to this reference measure, which leads to $\overline{q}$ dropping out entirely as presented in Proposition 2. Note that if $\mathbb{Q}\ll\overline{\mathbb{Q}}$, then $\overline{\mathbb{Q}}$ itself may serve as such a $\mathcal{G}$ invariant reference measure. We leverage this in Section A, where the Lebesgue measure serves as such a reference measure.
Proposition 2 (Invariant reference measure).
Suppose $\mathbb{H}$ is a $\mathcal{G}$ invariant unsigned measure that dominates both $\mathbb{Q}$ and $\overline{\mathbb{Q}}$. Then, $q/\overline{q}$ may be replaced with $q$ in Theorem 6, Theorem 7 and Corollaries 2 and 3.
Proof.
This follows immediately from the fact that it implies $\overline{q}$ is constant on each orbit, so that it is absorbed by the normalization constants in Theorem 7 and Corollaries 2 and 3, and by $c_{\alpha}^{[y]}$ in Theorem 6. ∎
Remark 9.
Theorem 6 implies that plugging the test statistic $T=q/\overline{q}$ in the traditional group invariance test yields the Neyman–Pearson-optimal test. In case we have densities with respect to a $\mathcal{G}$ invariant reference measure, then Proposition 2 reduces this to $T=q$, which is the main result in Lehmann and Stein (1949). Their proof strategy seems to implicitly rely on the assumption that $q$ is a density with respect to an available $\mathcal{G}$ invariant reference measure.
The log-optimal e-value can be viewed as the likelihood ratio between the alternative $\mathbb{Q}$ and its reverse information projection (RIPr) onto the null hypothesis (Larsson et al., 2024; Grünwald et al., 2024; Lardy et al., 2024). If $\mathbb{Q}\ll\overline{\mathbb{Q}}$, it turns out that the RIPr is $\overline{\mathbb{Q}}$. This condition holds, for example, if $\mathcal{G}$ is finite, as this means $\overline{\mathbb{Q}}$ is a finite average of measures, one of which is $\mathbb{Q}$. The proof is presented in Appendix E.8, and relies on Theorem 4.1 in Larsson et al. (2024).
Proposition 3 (RIPr).
Assume $\mathbb{Q}\ll\overline{\mathbb{Q}}$. Let $q$ denote the density of $\mathbb{Q}$ with respect to $\overline{\mathbb{Q}}$. Then $\overline{\mathbb{Q}}$ is a RIPr, and a log-optimal e-value is
$\varepsilon^{\textnormal{log}}=\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}=q.$
4.4 Expected utility-optimal e-values for group invariance: alternatives on orbits
In Section 4.3, we specified an aggregate objective function on the entire sample space $\mathcal{Y}$, and then applied Theorem 5 to decompose this into local objectives on each orbit. In this section, we use the insight in Remark 6 to turn Theorem 5 around: we specify an objective on each orbit separately, which then implies the resulting e-value is optimal uniformly in every aggregation of these orbit-level objectives.
To make this more concrete, we may specify a different alternative $\mathbb{Q}^{O}$ on each orbit $O\in\mathcal{Y}/\mathcal{G}$, and maximize some expected utility $\mathbb{E}^{\mathbb{Q}^{O}}[U(\varepsilon)]$ on each orbit. The composition $\varepsilon^{*}(y)=\varepsilon_{|O_{y}}^{*}(y)$ of the resulting orbit-level optimal e-values $\varepsilon_{|O_{y}}^{*}(y)$ is then expected utility-optimal for every marginal distribution over the orbits. While specifying an alternative $\mathbb{Q}^{O}$ on each orbit may sound like an arduous exercise, the following example shows its practical relevance. We put this example to practice in Section 7.3, where we apply it to the real-world experiment of Gilovich et al. (1985).
In practice, we only need to formulate the conditional alternative on the orbit that we actually observe. More precisely, when we receive our data $Y$, we may first classify its orbit $O_{Y}$, then choose an alternative $\mathbb{Q}^{O_{Y}}$ based on $O_{Y}$ (not based on $Y$ itself), and then compute the e-value using $Y$. Such an e-value is then optimal conditionally on the orbit, and also valid for group invariance. This does, of course, make it more difficult to explicitly articulate its unconditional optimality, as that would depend on the unobserved counterfactual of what we would have done in case we had observed other orbits. However, as this does not affect the validity, this may not be a problem in practice.
Example 9 (Hot hand).
The hot hand is a concept derived from basketball. It describes a momentum effect, in which a player hitting a shot increases their probability of hitting subsequent shots. This concept was first statistically popularized by Gilovich et al. (1985). In the recent literature, the hot hand is often examined by testing whether a sequence of shot outcomes $Y$ (hit/miss) is exchangeable against some sequential-dependence alternative that describes the hot hand effect (Miller and Sanjurjo, 2018; Ritzwoller and Romano, 2022).
Note that permuting such a shot sequence exactly fixes the statistic (#hit, #miss), that captures the number of hits (#hit) and the number of misses (#miss). This means the orbits under permutations may be labeled by (#hit, #miss).
If we were to follow the strategy of Section 4.3, we would be forced to specify a marginal alternative distribution over the number of hits and misses (over the orbits). This is hard in practice, as it requires knowledge of the skill of the player. In addition, if the data comes from actual basketball games it also requires knowledge of the strength of the opponent, as well as the teammates. This is further complicated by the fact that the hot hand phenomenon may, for example, be stronger for weaker players than highly skilled players.
The strategy in this section relieves us from specifying the marginal distribution over the orbits. That is, we need not specify the distribution of (#$(\#\textnormal{hit},\#\textnormal{miss})$: we merely need to specify the conditional distribution of the order of the hits and misses under the hot hand, given the statistic (#$(\#\textnormal{hit},\#\textnormal{miss})$. Here, we may even decide to use the number of hits and misses to influence the strength of the hot hand, as a proxy for skill. The resulting e-value is automatically optimal, uniformly in any distribution over (#$(\#\textnormal{hit},\#\textnormal{miss})$.
4.5 Optimal e-values for group invariance: objective on group
A final strategy is to specify an alternative $\mathbb{Q}^{\mathcal{G}}$ on the group $\mathcal{G}$ itself and test this against the Haar measure $\textnormal{Unif}(\mathcal{G})$ on the group. Such an e-value is valid for $\mathcal{G}$ invariance if and only if $\mathbb{E}_{\overline{G}}[\varepsilon(\overline{G})]\leq 1$. This may sound overly exotic, but it is not. In fact, this effectively underlies rank tests and sign tests, and thereby also almost all methods studied in the popular framework of conformal prediction.
To apply this idea, we may use an inversion kernel, as discussed in Section 2.7. In particular, we may derive an optimal e-value $\varepsilon_{\mathcal{G}}^{*}:\mathcal{G}\to[0,\infty]$ on the group for measuring evidence against the Haar measure, and then evaluate it using the inversion kernel $\gamma:\mathcal{Y}\to\mathcal{G}$: $\varepsilon_{\mathcal{G}}^{*}(\gamma(Y))$.
A concrete example is given by the rank statistic, which is in bijection with the group of permutations if we ignore ties. This means we may reimagine the function $\text{Rank}:\mathbb{R}^{M}\to\{1,\dots,M\}^{M}$ as a function that maps an observation on $\mathbb{R}^{M}$ to a group of permutations $\text{Rank}:\mathbb{R}^{M}\to\mathcal{G}$. Under exchangeability (permutation invariance), the distribution of the resulting rank is uniform (Haar) on the set of possible ranks. An alternative on the group corresponds to any other desired distribution on the ranks. If desired, such an alternative on the group may be obtained by pushing forward an alternative on $\mathcal{Y}$ to $\mathcal{G}$ through the inversion kernel.
5 E-processes for group invariance
5.1 Characterizing e-processes for group invariance
In this section, we describe sequential gathering of evidence against group invariance through e-processes. We define an e-process with respect to a hypothesis $H$ as a non-negative stochastic process $(\varepsilon_{n})_{n\geq 0}$, with $\varepsilon_{n}:\mathcal{X}\to[0,\infty]$, that is adapted to a filtration $(\mathcal{I}_{n})_{n\geq 0}$. We say an e-process is valid if
$\sup_{\mathbb{P}\in H}\mathbb{E}^{\mathbb{P}}[\varepsilon_{\tau}]\leq 1,$
for every stopping time $\tau$ adapted to the filtration $(\mathcal{I}_{n})_{n\geq 0}$. It is typical to impose $\varepsilon_{0}=1$, without loss of generality.
In Theorem 8, we characterize e-processes for $\mathcal{G}$ invariance. This result may be viewed as a sequential analogue of Theorem 1. Its proof is found in Appendix E.9.
Theorem 8 (E-processes for $\mathcal{G}$ invariance).
The stochastic process $(\varepsilon_{n})_{n\geq 0}$ adapted to $(\mathcal{I}_{n})_{n\geq 0}$, $\varepsilon_{n}:\mathcal{X}\to[0,\infty]$, is a valid e-process with respect to the filtration $(\mathcal{I}_{n})_{n\geq 0}$ for $\mathcal{G}$ invariance if and only if
$\sup_{\tau\in\mathcal{T}}\mathbb{E}_{\overline{G}}[\varepsilon_{\tau}(\overline{G}x)]\leq 1,$
for every $x\in\mathcal{X}$, where $\mathcal{T}$ is the collection of stopping times with respect to $(\mathcal{I}_{n})_{n\geq 0}$.
5.2 Constructing an e-process as an infimum over orbit-wise martingales
While Theorem 8 characterizes e-processes, it is not particularly instructive as to how to construct one. To construct an e-process, we may pass to an equivalent definition studied by Ramdas et al. (2022a), as a non-negative stochastic process that is almost surely bounded from above by a $\mathbb{P}$-non-negative martingale for every $\mathbb{P}\in H$. They show that any admissible e-process is of the form
$\varepsilon_{n}:=\textnormal{ess\,inf}_{\mathbb{P}\in H}\,\varepsilon_{n}^{\mathbb{P}},$
where $(\varepsilon_{n}^{\mathbb{P}})_{n\geq 0}$ is a $\mathbb{P}$-non-negative martingale with respect to the same filtration that starts at 1.
In case of a large non-parametric hypothesis such as group invariance, taking the infimum over every $\mathcal{G}$ invariant probability may seem like an arduous exercise. Luckily, we find that it suffices to take the infimum over martingales for uniform distributions on orbits:
$\varepsilon_{n}=\textnormal{ess\,inf}_{O\in\mathcal{X}/\mathcal{G}}\,\varepsilon_{n}^{O},$
where $(\varepsilon_{n}^{O})_{n\geq 0}$, $\varepsilon_{n}^{O}:\mathcal{X}\to[0,\infty]$, is a martingale under $\textnormal{Unif}(O)$. We present this result in Theorem 9.
Note that being a martingale under $\textnormal{Unif}(O)$ does not restrict the behavior of $(\varepsilon_{n}^{O})_{n\geq 0}$ for $x\not\in O$. For this reason, we impose $\varepsilon_{n}^{O}(x)=\infty$ for $x\not\in O$ to ensure values outside the orbit do not affect the infimum. Indeed, it implies $\textnormal{ess\,inf}_{O\in\mathcal{X}/\mathcal{G}}\,\varepsilon_{n}^{O}(x)=\varepsilon_{n}^{O}(x)$ for $x\in O$, as the orbits partition the sample space. Its proof is presented in Appendix E.10.
Theorem 9 (E-process as infimum of orbit-based martingales).
Suppose that, for every $O\in\mathcal{X}/\mathcal{G}$, $(\varepsilon_{n}^{O})_{n\geq 0}$ is a non-negative supermartingale for $\textnormal{Unif}(O)$ adapted to the filtration $(\mathcal{I}_{n})_{n\geq 0}$ that starts at 1, and $\varepsilon_{n}^{O}(x)=\infty$ for $x\not\in O$. Then, $(\varepsilon_{n})_{n\geq 0}$ defined by
$\varepsilon_{n}=\textnormal{ess\,inf}_{O\in\mathcal{X}/\mathcal{G}}\,\varepsilon_{n}^{O}$
is an anytime valid e-process for $\mathcal{G}$ invariance with respect to $(\mathcal{I}_{n})_{n\geq 0}$.
Remark 10 (Comments on Theorem 9).
If the orbit $O$ of $X$ was known from the start (i.e. there is only one orbit), then taking the infimum is superfluous: we could immediately identify the orbit in which $X$ lands, and therefore simply run a martingale for $\textnormal{Unif}(O)$. In practice, the orbit of $X$ is often not known, but we may sequentially learn about it. The process described in Theorem 9 may be viewed as tracking a martingale for each candidate orbit that $X$ may eventually turn out to be a member of, and ‘dropping’ martingales as soon as their associated orbit is no longer a candidate by setting them to $\infty$ so that they do not affect the infimum.
5.3 Constructing orbit-wise martingales out of an e-value for $\mathcal{G}$ invariance
A simple way to construct the orbit-wise martingales for Theorem 9 is given in Proposition 4, which is inspired by Theorem 2 in Koning and van Meer (2025). The idea is to start by picking some information horizon $\mathcal{I}$, where $\mathcal{I}\supseteq\mathcal{I}_{n}$, $n\geq 0$. At this horizon $\mathcal{I}$, we formulate some desired $\mathcal{I}$-measurable e-values $\varepsilon^{O}$. Then, we induce (Doob) martingales for each $\varepsilon^{O}$ and track their running infimum.
This strategy nicely harmonizes with our discussion on optimal e-values in Section 4. There, the idea was to establish an optimal e-value on each orbit, and then aggregate these. To apply the outlined strategy, we may reverse this process: we derive some optimal e-value $\varepsilon^{*}$, and then retrieve orbit-level e-values $\varepsilon_{|O}^{*}$ by restricting $\varepsilon^{*}$ to the orbits:
$\varepsilon_{|O}^{*}(x):=\begin{cases}\varepsilon^{*}(x)&\textnormal{ if }x\in O,\\ \infty,&\textnormal{ if }x\not\in O.\end{cases}$ (8)
These orbit-level e-values may then be used in the machinery of Proposition 4. Its proof is given in Appendix E.11.
The remarkable feature of this construction is that it results in an e-process that equals $\varepsilon^{*}$ if the information horizon $\mathcal{I}$ is reached. We illustrate this process in Example 10, where we construct a (non-martingale) e-process in this manner.
Proposition 4 (Inducing orbit-wise martingales).
Let $\varepsilon^{*}:\mathcal{X}\to[0,\infty]$ be a valid $\mathcal{I}$-measurable e-value for $\mathcal{G}$ invariance, $\mathcal{I}\supseteq\mathcal{I}_{n}$, $n\geq 0$. For every orbit $O\in\mathcal{X}/\mathcal{G}$, define
$\varepsilon_{n}^{O}=\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{|O}^{*}\mid\mathcal{I}_{n}].$
Then, $(\varepsilon_{n}^{O})_{n\geq 0}$ is a martingale for $\textnormal{Unif}(O)$ with respect to $(\mathcal{I}_{n})_{n\geq 0}$ that starts at 1. Hence,
$\varepsilon_{n}=\textnormal{ess\,inf}_{O\in\mathcal{X}/\mathcal{G}}\,\varepsilon_{n}^{O},$
is a valid e-process for $\mathcal{G}$ invariance. If $\mathcal{I}_{N}=\mathcal{I}$ for some $N\geq 0$, then $\varepsilon_{N}=\varepsilon^{*}$.
Example 10 (Illustration of e-process).
We discuss a simple example of a non-trivial e-process, to illustrate the results in this section.
We are to sequentially observe a pair of letters, $\mathcal{X}=\{AB,BA,AC,CA\}$. Let $\mathcal{G}$ be the group that permutes the two letters. This means we have two orbits: $O_{1}=\{AB,BA\}$ and $O_{2}=\{AC,CA\}$. Let $X$ denote our random variable on $\mathcal{X}$, and let $X^{1}$ denote the first letter of $X$ and $X^{2}=X$. Let $\mathcal{I}_{1}=\sigma(X^{1})$ and $\mathcal{I}_{2}=\sigma(X^{2})=\sigma(X)$, so that we sequentially observe two letters.
As an example, we consider the log-optimal e-value against any aggregation of the orbit-wise alternatives $\mathbb{Q}_{1}(AB)=2/3$, $\mathbb{Q}_{1}(BA)=1/3$ and $\mathbb{Q}_{2}(AC)=1/3$, $\mathbb{Q}_{2}(CA)=2/3$. By Theorem 5, this is given by
$\varepsilon_{2}(AB)=4/3,\quad\varepsilon_{2}(BA)=2/3,\quad\varepsilon_{2}(AC)=2/3,\quad\varepsilon_{2}(CA)=4/3,$
which automatically gives the value of the e-process at time 2.
To obtain the e-process at time 1, we now restrict these e-values to the orbits as in (8), and apply the construction in Proposition 4 to obtain
$\varepsilon_{1}^{O_{1}}(A)$ $=4/3,$ $\varepsilon_{1}^{O_{1}}(B)$ $=2/3,$ $\varepsilon_{1}^{O_{1}}(C)$ ==$\infty$,= ∞ ,
$\varepsilon_{1}^{O_{2}}(A)$ $=2/3,$ $\varepsilon_{1}^{O_{2}}(B)$ ==$\infty$,= ∞ , $\varepsilon_{1}^{O_{2}}(C)$ $=4/3.$
Minimizing over the two orbits yields the e-process at time 1:
$\varepsilon_{1}(A)=2/3,\quad\varepsilon_{1}(B)=2/3,\quad\varepsilon_{1}(C)=4/3.$
It is straightforward, though somewhat tedious, to check that this process indeed has expectation bounded by 1 for all 8 possible stopping times under uniformity on each orbit. We also stress that this e-process is not a supermartingale for $\mathcal{G}$ invariance, as
$\sup_{\mathbb{P}:\mathcal{G}\textnormal{-invariant}}\mathbb{E}^{\mathbb{P}}[\varepsilon_{2}\mid X^{1}=A]\geq\max_{i}\mathbb{E}^{\textnormal{Unif}(O_{i})}[\varepsilon_{2}\mid X^{1}=A]=\max_{i\in\{1,2\}}\varepsilon_{1}^{O_{i}}(A)=4/3>2/3=\varepsilon_{1}(A).$
Remark 11 (Non-compact groups and de Finetti).
As in the non-sequential setting discussed in Remark 4, we may generalize the characterization of e-processes to non-compact groups through an ergodic theorem. Here, the uniform probabilities on orbits are replaced by ergodic probabilities. To construct an e-process in such a setting, we may then track an infimum of martingales for the ergodic probabilities, instead of tracking an infimum over martingales for each $\mathcal{G}$ invariant probability.
Of course, without imposing additional structure, tracking a martingale for each ergodic probability may still be a daunting task. To highlight this, we may consider perhaps the most famous example of an ergodic theorem: de Finetti’s theorem. Under suitable regularity conditions, it states that if an infinite sequence is exchangeable (invariant under permutations that move finitely many elements), then its law may be written as a mixture over i.i.d. probabilities (the ergodic measures). This has been explored by Ramdas et al. (2022b) in a binary and $d$-ary setting, to show the existence of e-processes in settings where no powerful martingales exist. Unfortunately, tracking a martingale for each i.i.d. probability is practically difficult beyond simple examples such as binary data.
6 Test martingales for group invariance
One of the popular features of traditional group invariance tests is that they are easy to operationalize; a property we find extends to e-values in Section 3. Indeed, by observing the data $X=x$, we simultaneously observe the orbit $O_{x}$, and we may easily sample from this orbit through $\overline{G}x$. Unfortunately, in the sequential context discussed in Section 5, this benefit breaks down as we need to keep track of a martingale for each orbit, which may be impractical.
In this section, we consider test martingales for group invariance as a practical alternative. A test martingale for a hypothesis $H$ is defined as a non-negative stochastic process $(\varepsilon_{n})_{n\geq 0}$ adapted to a filtration $(\mathcal{I}_{n})_{n\geq 0}$ that starts at 1 and satisfies
$\sup_{\mathbb{P}\in H}\mathbb{E}^{\mathbb{P}}[\varepsilon_{n+1}\mid\mathcal{I}_{n}]\leq\varepsilon_{n},$
for every $n\geq 0$. Such test martingales are also e-processes.
In this section, we show how we may build such a test martingale for group invariance under a natural filtration. In particular, we consider a filtration that matches the group setting, sequentially revealing a growing sequence of subgroups $(\mathcal{G}_{n})_{n\geq 0}$ and data $(X^{n})_{n\geq 0}$. At each point in time $n$, we show how to construct an e-value for $\mathcal{G}_{n}$ invariance of $X^{n}$ conditional on the past data $X^{n-1}$, and then construct a martingale as their sequential product.
In this setup we actually need not pre-specify the group structure, which may have practical advantages. For example, an analyst may sequentially receive batches of data of different size, according to some unknown process. In such a setting, the analyst may only know what invariance they want to test when (the metadata of) the new batch has arrived; see Example 12.
An under-appreciated fact about test martingales is that they offer a stronger guarantee than e-processes: they offer validity even if the data is generated by a different distribution in the null hypothesis at each moment in time (see e.g. the notion of fork-convexity in Ramdas et al. (2022b)). The downside of test martingales is that this stronger guarantee makes them less powerful than e-processes in settings where this additional guarantee is not important; see Remark 12 and the examples that follow it for a discussion.
In Appendix B, we discuss how we may reduce to a single orbit. This yields a simple null hypothesis: a uniform distribution on this orbit, so that admissible e-processes and martingales coincide.
6.1 A sequence of random variables invariant under a sequence of groups
To introduce this setting, we first present some notation and concepts. We embed the entire sequential setting in a latent sample space $\mathcal{X}$. In particular, we assume we have a nested sequence of subspaces $(\mathcal{X}^{n})_{n\geq 0}$ of $\mathcal{X}$: $\mathcal{X}^{n}\subseteq\mathcal{X}^{n+1}$, which are tied together through a sequence of continuous maps $(\textnormal{proj}_{\mathcal{X}^{n}})_{n\geq 0}$ which project onto the subsets, $\textnormal{proj}_{\mathcal{X}^{n}}:\mathcal{X}\to\mathcal{X}^{n}$. With a projection map, we mean that such a map satisfies $\textnormal{proj}_{\mathcal{X}^{n}}(x)=x$ if $x\in\mathcal{X}^{n}$.
[…]
Our goal now is to test the hypothesis that $(X^{n})_{n\geq 0}$ is invariant under $(\mathcal{G}_{n})_{n\geq 0}$.
Example 11 (Exchangeability and i.i.d.).
Suppose that $X^{n}=(Y_{0},\dots,Y_{n})$ for each $n$. Let us choose $\mathcal{G}_{n}=\mathfrak{P}_{n}$ as the group of permutations on $n+1$ elements. If $(X^{n})_{n\geq 0}$ is invariant under $(\mathfrak{P}_{n})_{n\geq 0}$, then we say that $(X^{n})_{n\geq 0}$ is exchangeable. Referring back to Remark 11, testing exchangeability is equivalent to testing whether the sequence is i.i.d. by de Finetti’s theorem.
Example 12 (Within-batch exchangeability).
Suppose we sequentially observe potentially unequally sized batches of data $Y_{0},Y_{1},\dots$, where each $Y_{i}$ is exchangeable, $i=0,1,\dots$. We can choose $\mathcal{G}_{n}=\mathfrak{P}^{0}\times\mathfrak{P}^{1}\times\cdots\times\mathfrak{P}^{n}$, where $\mathfrak{P}^{i}$ is the group of permutations acting on the batch $Y_{i}$. Defining $X^{n}=(Y_{0},\dots Y_{n})$, within-batch exchangeability can be viewed as invariance of $(X^{n})_{n\geq 0}$ under this group $(\mathcal{G}_{n})_{n\geq 0}$.
If we view the elements of a batch as individual observations, then within-batch exchangeability is weaker than exchangeability of individual observations: we exclude permutations that swap observations across batches. Specifically, the groups we consider here are subgroups of the permutations on the set of the individual observations. The idea to test sequential invariance of all observations by batching units into pairs has been independently explored by Saha and Ramdas (2024).
6.2 Filtration
We are now ready to discuss the filtration. We assume that at each moment in time, we know the past data $X^{n-1}$ as well as the group $\mathcal{G}_{n}$ under which we are about to test invariance. This means that $X^{n}$ should be $\mathcal{I}_{n}$-measurable and $\mathcal{G}_{n}$ is $\mathcal{I}_{n-1}$-measurable (‘predictable’), which leads to a filtration of the form
$\sigma(\mathcal{G}_{0})\subseteq\sigma(\mathcal{G}_{1},X^{0})\subseteq\sigma(\mathcal{G}_{2},X^{1})\subseteq\sigma(\mathcal{G}_{3},X^{2})\subseteq\cdots.$
To construct our martingale, we will formulate ‘conditional’ e-values, which are valid for $\mathcal{G}_{n}$ invariance, conditional on the past information. Recall from Section 3.2 that testing invariance is equivalent to testing uniformity conditionally on the drawn orbit. By conditioning on the orbit, we are effectively squeezing an additional step into the flow of information which reveals the orbit $O_{X^{n}}$ before it reveals $X^{n}$ itself:
$\sigma(\mathcal{G}_{0})\subseteq\sigma(\mathcal{G}_{0},O_{X^{0}})\subseteq\sigma(\mathcal{G}_{1},X^{0})\subseteq\sigma(\mathcal{G}_{1},O_{X^{1}},X^{0})\subseteq\sigma(\mathcal{G}_{2},X^{1})\subseteq\cdots.$
Luckily we may compress this unwieldy filtration. First, in the context of testing invariance, learning the group or orbit is not informative for whether our data is invariant under this group / on this orbit. This means we may compress time steps at which the group or orbit are revealed:
$\sigma(\mathcal{G}_{0},O_{X^{0}})\subseteq\sigma(\mathcal{G}_{1},O_{X^{1}},X^{0})\subseteq\sigma(\mathcal{G}_{2},O_{X^{2}},X^{1})\subseteq\sigma(\mathcal{G}_{3},O_{X^{3}},X^{2})\subseteq\cdots.$
Second, once the orbit is determined, then the precise group that brought us to this orbit is irrelevant (different groups may induce the same orbit). For this reason, we consider the compressed filtration that describes the flow of information as
$\sigma(O_{X^{0}})\subseteq\sigma(X^{0},O_{X^{1}})\subseteq\sigma(X^{1},O_{X^{2}})\subseteq\sigma(X^{2},O_{X^{3}})\subseteq\cdots.$
We use the shorthand $\mathcal{I}_{n}=\sigma(X^{n},O_{X^{n+1}})$, $n\geq 0$ and $\mathcal{I}_{0}=\sigma(O_{X^{0}})$.
6.3 Test martingale
To construct a test martingale, we construct conditional e-values $\varepsilon_{n}$, that are valid for $\mathcal{G}_{n}$ invariance conditional on the past data:
$\mathbb{E}^{\textnormal{Unif}(O_{X^{n}})}[\varepsilon_{n}\mid\mathcal{I}_{n-1}]\leq 1.$ (9)
The test martingale itself is then given by its running product:
$\varepsilon^{n}(X^{n})=\prod_{i=0}^{n}\varepsilon_{n}(X^{n}).$
In Proposition 5, we present a characterization of the conditional e-value (9). The trick underlying this result is captured in Lemma 2, which characterizes the conditional distribution $\textnormal{Unif}(O_{X^{n}})\mid(X^{n-1},O_{X^{n}})$ by means of a subgroup that stabilizes the past data. In particular, given $X^{n-1}=x^{n-1}$, we define $\mathcal{K}_{n}(x^{n-1})=\{G\in\mathcal{G}_{n}:Gx^{n-1}=x^{n-1}\}$, for $n\geq 1$ and $\mathcal{K}_{0}=\mathcal{G}_{0}$. In Section E.12 in the Supplementary Material, we show that this is indeed a compact subgroup of $\mathcal{G}_{n}$, and include a proof of a more general result.
Proposition 5.
An e-value $\varepsilon_{n}$ is conditionally valid given $O_{X^{n}}$ and $X^{n-1}$ when $\mathbb{E}_{\overline{K}_{n}}[\varepsilon_{n}(\overline{K}_{n}x^{n})]\leq 1,\textnormal{ for every }x^{n}\in\mathcal{X}^{n}$.
Lemma 2.
[…]
Proposition 5 shows that we may reduce the problem of constructing a conditional e-value to constructing an unconditional e-value that is valid for invariance under a data-dependent group $\mathcal{K}_{n}(X^{n-1})$. This means we may immediately apply the machinery derived in Section 3 and 4, where we study the construction of such unconditional e-values. For example, following Section 3, we may choose
$\varepsilon_{n}(X^{n})=\frac{T_{n}(X^{n})}{\mathbb{E}_{\overline{K}_{n}}T_{n}(\overline{K}_{n}X^{n})},$
where $T_{n}$ is a predictable non-negative test statistic.
Alternatively, given a predictable alternative $\mathbb{Q}_{n}$ on $\mathcal{X}^{n}$, we may define ¯$\overline{\mathbb{Q}}_{n}=\mathbb{E}_{\overline{K}_{n}}[\overline{K}_{n}\mathbb{Q}_{n}]$ with densities $q_{n}$ and $\overline{q}_{n}$ with respect to some reference measure and construct an expected utility-optimal e-value as in Section 4 of the form
$\varepsilon_{n}(X^{n})=(U^{\prime})^{-1}\left(\lambda^{*}\frac{\overline{q}_{n}(X^{n})}{q_{n}(X^{n})}\right).$
Remark 12.
When using such a test martingale, we are effectively testing whether $(X^{n})_{n\geq 1}$ is $(\mathcal{K}_{n})_{n\geq 0}$ invariant. As the subgroups $(\mathcal{K}_{n})_{n\geq 0}$ may be less rich than the original groups, we are testing a larger null hypothesis than $(\mathcal{G}_{n})_{n\geq 0}$ invariance. This reveals where the test martingale loses power compared to an e-process. We illustrate this in Example 13, 14 and 15.
Example 13 (Sequential sphericity).
Suppose that $\mathcal{X}^{n}=\mathbb{R}^{n}$ so that $X^{n}$ is a random $n$-vector for all $n$. Let $\mathcal{O}_{n}$ be the collection of $n\times n$ orthonormal matrices. Then, $X^{n}$ is said to be spherically distributed if it is invariant under $\mathcal{O}_{n}$. We consider testing invariance of the sequence $(X^{n})_{n\geq 1}$ under matrix multiplication by the orthonormal matrices in $(\mathcal{O}_{n})_{n\geq 1}$.
In this example, the orbit $O_{X^{n}}$ is the hypersphere in $n$ dimensions that contains $X^{n}$. As a consequence, the effective filtration reveals the previous observations $X^{n-1}$ and the length of $X^{n}$. Together, these determine $X^{n}$ up to the sign of its final element. As a result, $\mathcal{K}_{n}$ contains two elements: $\textnormal{diag}(1,\dots,1,1)$ and $\textnormal{diag}(1,\dots,1,-1)$, which flips the sign of the final element. This is equivalent to testing whether $X^{n}$ is invariant under sign-flips.
Example 14 (Test martingale for exchangeability).
Continuing from Example 11, suppose we sequentially observe $X^{n}=(Y_{0},Y_{1},\dots,Y_{n})$ that are exchangeable.
Here, it turns out that $X^{n}$ is degenerate conditional on $\sigma(X^{n-1},O_{X^{n}})$. In particular, $X^{n-1}=(Y_{0},Y_{1},\dots,Y_{n-1})$ and $O_{X^{n}}$ equals the multiset $\{Y_{0},\dots Y_{n}\}$. Hence, $Y_{n}$ is simply the value in $O_{X^{n}}$ that is not accounted for in $X^{n-1}$. As a consequence, the conditional distribution $X^{n}$ given $X^{n-1}$ and $O_{X^{n}}$ is degenerate. Assuming the realizations are distinct, this means $\mathcal{K}_{n}$ only contains the identity element for each $n$.
A consequence is that it is impossible to sequentially test exchangeability with a test martingale under the filtration $(\sigma(X^{n}))_{n\geq 0}$, as previously observed by Vovk (2021) and Ramdas et al. (2022b). Our discussion gives some context around their impossibility result, by showing it may be interpreted as the group $\mathcal{K}_{n}$ becoming degenerate.
Example 15 (Test martingale for within-batch exchangeability).
Continuing from Example 12, let us again consider $X^{n}=(Y_{0},\dots Y_{n})$, where each $Y_{i}$ is an exchangeable batch of data. Let us assume the realizations are distinct in each batch. Then, $\mathcal{K}_{n}(X^{n-1})=\{I^{0}\}\times\{I^{1}\}\times\cdots\times\{I^{n-1}\}\times\mathfrak{P}^{n}$, where $I^{i}$ denotes the identity permutation acting on the $i$th batch, for $n\geq 1$ and $\mathcal{K}_{0}=\mathfrak{P}^{0}$. That is, the conditional distribution of $X^{n}$ is uniform on the final batch. Interestingly, the stabilizer $\mathcal{K}_{n}(X^{n-1})$ does not depend on $X^{n-1}$.
As discussed in Example 12, exchangeability implies within-batch exchangeability. This means rejecting within-batch exchangeability also rejects exchangeability. As a result, we can construct a sequential test for exchangeability by merging observations into batches. This of course impoverishes the filtration, since we only look at the data after a batch has arrived. The size of a batch is allowed to be adaptive. Generalizing this reasoning is the topic of Appendix B.
7 Simulations and application
7.1 Case-control experiment and learning the alternative
In this simulation study, we consider a hypothetical case-control experiment in which units are assigned to either the treated or control set uniformly at random. In each interval of time, we receive the outcomes of a number of treated and control units, where the number of treated and control units is Poisson distributed with parameter $\theta>0$ and a minimum of 1. The outcomes of the treated units are $\mathcal{N}(a,1)$-distributed and the outcomes of the controls are $\mathcal{N}(b,1)$-distributed. The true mean and variance are considered unknown, and are adaptively learned based on the previously arrived data. As a batch of data, we consider the combined observations of both the treated and control units that arrived in the previous interval of time.
As a result, a batch $X_{t}$ of $n^{t}$ outcomes, consisting of $n_{a}^{t}$ treated and $n_{b}^{t}$ control units, can be represented as
$X_{t}\sim\begin{bmatrix}1_{n_{a}^{t}}a\\ 1_{n_{b}^{t}}b\end{bmatrix}+\mathcal{N}\left(0,I\right),$
where $1_{n_{a}^{t}}$ and $1_{n_{b}^{t}}$ denote vectors of $n_{a}^{t}$ and $n_{b}^{t}$ ones, respectively, and the first $n_{a}^{t}$ elements correspond to the treated units, without loss of generality. We would like to base our test statistic on the difference of sample means: $\overline{1}_{n^{t}}^{\prime}X_{t}\sim\mathcal{N}\left(a-b,1/n_{a}^{t}+1/n_{b}^{t}\right)$, where $\overline{1}_{n^{t}}=(1_{n_{a}^{t}}(n_{a}^{t})^{-1},-1_{n_{b}^{t}}(n_{b}^{t})^{-1})$. In particular, we will test the null hypothesis that the elements of a batch $X_{t}$ are exchangeable and so $a=b$, against the alternative hypothesis that $a>b$.
We use a test-martingale-based on the log-optimal e-value for testing exchangeability against our current estimate of the Gaussian alternative, as derived in Appendix A,
$\varepsilon_{t}=\frac{\exp\{(\widehat{a}_{t-1}-\widehat{b}_{t-1})/\widehat{\sigma}_{t-1}^{2}\times\overline{1}_{n^{t}}^{\prime}X_{t}\}}{\mathbb{E}_{\overline{G}}\exp\{(\widehat{a}_{t-1}-\widehat{b}_{t-1})/\widehat{\sigma}_{t-1}^{2}\times\overline{1}_{n^{t}}^{\prime}\overline{G}X_{t}\}},$
where $\widehat{a}_{t-1}-\widehat{b}_{t-1}=\overline{1}_{n^{t-1}}^{\prime}X_{t-1}$ is our treatment estimator at time $t-1$ and $\widehat{\sigma}_{t-1}^{2}$ is its pooled sample variance estimator, and $\overline{G}$ is uniform on the permutations of $n^{t}$ elements. For the first batch, we can either rely on an educated guess, or skip it for inference and only use it for estimating these parameters. We estimate the normalization constant by using 100 permutations drawn uniformly at random with replacement.
[…]
In Figure 1, we plot the test-martingale-based e-processes for 1 000 simulations. The dotted line indicates the value $20=1/0.05$, so that exceeding this line corresponds to a rejection at level $\alpha=0.05$. The plot on the left features the setting under the null, and the plot on the right the setting under the alternative. To make the figure easier to interpret, we plot at each time the line below which 5%, 50% and 95% of the test martingales have remained up until that point. For example, in the right plot, roughly 95% of the e-processes have exceeded 20 at batch 23, so that the power at level $\alpha$ is roughly 95% after 23 batches. As expected, the left plot shows that 95% of the e-processes remain below 20 under the null.
Figure 1: Plots of 1 000 e-processes over the number of arrived batches. The highlighted lines are running quantiles: xof the e-processes have not crossed above the line at the indicated time. The plot on the left is under the null hypothesis, and the plot on the right is under the alternative. The horizontal dotted line is at 20.
7.2 Testing symmetry and comparison to de la Peña (1999)
In this simulation study, we consider testing sign-symmetry of data as in Appendix A.7. We compare our e-process to the one based on de la Peña (1999) when testing against a simple normal alternative $X_{i}\sim\mathcal{N}(m,1)$ with $m=1$.
We plot 1 000 e-processes of each type in Figure 2. The plot on the left is our log-optimal e-value-based e-process, whereas the plot on the right is based on de la Peña (1999). The figure shows that our e-processes grow much more quickly. This coincides with the observation made by Ramdas et al. (2022a) that the e-process based on de la Peña (1999) is inadmissible.
Figure 2: Plots of 1 000 e-processes over the number of arrived observations under a normal alternative with mean $m=1$. The highlighted lines are running quantiles: xof the e-processes have not crossed above the line at the indicated time. The plot on the left is for our log-optimal e-value-based e-process, and the plot on the right is for the one based on de la Peña (1999). The horizontal dotted line is at 20.
7.3 Illustrative application: optimal e-values for the hot hand
Following Remark 9, we illustrate our methodology by applying it to testing the hot hand in basketball. [^7] The hot hand methodology described here was developed during the supervision of many master’s thesis projects on this topic, including the theses of Bette Donker, Junda Fu, Max Broers, Nidas Brandsma, Nicole Serban, Raslen Kouzana, Sam Hammink, Stijn Koene, Tijn Wouters and Lloyd Vissers, who applied this methodology and preceding variations to a variety of sports and settings. Suppose we are to observe the outcomes (hits/misses) of $n$ shots of a basketball player. The hot hand hypothesis is that there exists some sequential dependence in these outcomes. In particular, if a player hits a number of shots then they are hypothesized to be more successful in hitting subsequent shots. To test the hot hand, it is common to test the null hypothesis of exchangeability, which bans any sequential dependence (Miller and Sanjurjo, 2018; Ritzwoller and Romano, 2022).
To specify an e-value, we may condition on (#hit, #miss), which is equivalent to conditioning on the orbit. This orbit contains each order of hits and misses for a given value of (#hit, #miss). Under exchangeability, each arrangement of hits and misses on the orbit is equally likely. Following Section 4.4, we may also specify the alternative conditional on the orbit, to describe how we believe the hot hand works. This absolves us from having to specify an alternative across orbits, which would require prior knowledge of the skills of the shooters. The resulting e-value is automatically optimal uniformly over mixtures across orbits by Theorem 5.
For simplicity, let us assume that a player is ‘hot’ if they hit $k$ shots in a row. If a player is hot, suppose this boosts their probability of hitting the next shot conditional on the orbit through $p_{\textnormal{hot}}=(p_{\textnormal{neutral}})^{\beta}$, where $p_{\textnormal{neutral}}$ represents the conditional probability to hit in the absence of a hot hand: the number of remaining hits divided by the number of remaining shots in the sequence. This means that if $\beta=1$, $p_{\textnormal{hot}}=p_{\textnormal{neutral}}$, and $p_{\textnormal{hot}}>p_{\textnormal{neutral}}$ when $\beta<1$. For example, if $p_{\textnormal{neutral}}=0.5$ and $\beta=0.9$, then $p_{\textnormal{hot}}\approx 0.536$ — a modest boost.
For example, suppose the shot sequence is 111011, where 1 represents a hit and 0 a miss, and let $\beta=0.9$ and $k=2$. As there are $\binom{6}{4}=15$ permutations of this sequence, the conditional probability of this sequence given the orbit equals 1/15 under the null. Under the alternative, we decompose the conditional probability of the shot sequence given (#hit, #miss), into a sequence of further conditional probabilities given the previous shot outcomes:
$\textnormal{Pr}_{\beta=0.9}(111011\mid(4,2))$
$=\textnormal{Pr}_{\beta=0.9}(\textnormal{shot }1=1\mid(4,2))\times\textnormal{Pr}_{\beta=0.9}(\textnormal{shot }2=1\mid(4,2),\textnormal{shot }1=1)\times\cdots$
$=4/6\times 3/5\times(2/4)^{0.9}\times(1-(1/3)^{0.9})\times 1/2\times 1\approx 0.0673.$
where the powers of $0.9$ are because the preceding two shots were a hit, increasing the probability of a subsequent hit. By Corollary 2, the resulting log-optimal e-value equals $\approx 0.0673/(1/15)=1.0095$ — tiny evidence against no hot hand. Such a tiny bit of evidence is not unexpected in the context of the hot hand. Indeed, in the context of traditional permutation testing, Ritzwoller and Romano (2022) argue that an individual binary shot sequence contains little information to discriminate between the hypotheses so that we need long sequences to potentially detect a hot hand.
We offer an alternative solution to the problem that an individual shot sequence does not contain much evidence, by using a powerful merging property of e-values: the product of independent e-values is also an e-value. This means we may compute an e-value for a large set of independent shot sequences, and aggregate their evidence by merging these e-values into a much more powerful e-value through multiplication.
We apply this idea to the controlled shooting experiment data collected by Gilovich et al. (1985), with 26 shooters taking up to 100 shots each. [^8] We retrieved this data from the Supplementary Material of Miller and Sanjurjo (2018). We consider variations of the hot hand that trigger after 1, 2 or 3 consecutive hits, with $\beta\in\{0.85,0.9\}$. Table 1 reports the product of the e-values for the individual shooters. To interpret these e-values, recall that its reciprocal $\mathfrak{p}=1/e$ is a post-hoc p-value, which we may interpret as a rejection at level $\mathfrak{p}$ under a generalized Type I error (Koning, 2024a; Grünwald, 2024). Looking at the product, we find strong evidence in support of the null (no hot hand) when compared to a 1-hit hot hand, but we find substantial evidence against the null for 2-hit and 3-hit triggers. This suggests the hot hand in such shooting experiments is only triggered after more than one hit. The full table with e-values for each shooter is reported in Appendix F.
Trigger 1 hit 2 hits 3 hits
$\beta$ 0.85 0.90 0.85 0.90 0.85 0.90
Product e-value 0.007 0.180 3.108 4.460 7.489 5.525
Post-hoc p-value 142.9 5.556 0.322 0.224 0.134 0.181
Table 1: Product of log-optimal e-values and post-hoc p-values ($\mathfrak{p}=1/e$) for the controlled shooting experiment of Gilovich et al. (1985) for exchangeability against several hot hand alternatives, triggering after 1-3 hits for a modest effect $(\beta=0.85)$ and weak effect $(\beta=0.9)$.
8 Acknowledgements
We thank Sam van Meer, Muriel Pérez-Ortiz, Tyron Lardy, Will Hartog, Yaniv Romano, Jake Soloff, Stan Koobs, Peter Grünwald, Wouter Koolen, Aaditya Ramdas and three anonymous referees for useful discussions and comments. Part of this research was performed while the author was visiting the Institute for Mathematical and Statistical Innovation (IMSI), which is supported by the National Science Foundation (Grant No. DMS-1929348).
Appendix A Illustration: optimal e-values for invariance against Gaussian location-shift
In this section, we illustrate our optimal e-values for testing invariance under a group of orthonormal matrices, against a Gaussian alternative under a location shift. If we include all orthonormal matrices, this yields clean connections to parametric theory and Student’s $t$-test. Moreover, we also consider exchangeability, which reveals an interesting relationship to the softmax function. In addition, we consider sign-symmetry, which we relate to a previously-studied e-value based on de la Peña [1999], and to work of Vovk and Wang [2024].
We start with an exposition of the invariance-based concepts for the orthogonal group $\mathcal{O}(d)$ that consists of all orthonormal matrices.
A.1 Sphericity
Suppose that $\mathcal{Y}=\mathbb{R}^{d}\setminus\{0\}$ and $\mathcal{G}=\mathcal{O}(d)$ is the orthogonal group, which can be represented as the collection of all $d\times d$ orthonormal matrices, $d\geq 1$. The orbits $O_{y}=\{z\in\mathcal{Y}\ |\ z=Gy,\exists G\in\mathcal{G}\}$ of $\mathcal{G}$ in $\mathbb{R}^{d}$ are the concentric $d$-dimensional hyperspheres about the origin. Each of these hyperspheres can be uniquely identified with their radius $\mu>0$. To obtain a $\mathcal{Y}$-valued orbit representative, we multiply $\mu$ by an arbitrary unit $d$-vector $\iota$ to obtain $\mu\iota$. For example $y$ lies on the orbit $O_{y}$ that is the $d$-dimensional hypersphere with radius $\|y\|_{2}$, and has orbit representative $[y]=\|y\|_{2}\iota$.
For simplicity, we now first focus on the subgroup $SO(2)$ of $O(2)$ and its action on $\mathbb{R}^{2}\setminus\{0\}$, which exactly describes the (orientation-preserving) rotations of the circle, and has the same orbits as $O(2)$. The reason we focus on $SO(2)$, is because its group acts freely on each concentric circle. As a consequence, every element in the group can be uniquely identified with an element on the unit circle $S^{1}$ (and in fact on every orbit). We choose to identify the identity element with $\iota$, and we identify every element of $SO(2)$ with the element on the circle that we obtain if that rotation is applied to $\iota$. We denote this induced group action of the unit circle $S^{1}$ on $\mathcal{Y}$ by $\circ$.
Under this bijection between the group and the unit circle, we can define our inversion kernel map $\gamma$ as $\gamma(y)=y/\|y\|_{2}$, which may be viewed as the group element that rotates $\iota$ to $y/\|y\|_{2}$. To see that $\gamma$ is indeed an inversion kernel, observe that
$\gamma(y)[y]=y/\|y\|_{2}\circ\iota\|y\|_{2}=[(y/\|y\|_{2})\circ\iota]\|y\|_{2}=(y/\|y\|_{2})\|y\|_{2}=y,$ (10)
where the second equality follows from the fact that the action of $(y/\|y\|_{2})$ on $\iota$, rotates $\iota$ to $y/\|y\|_{2}$. Invariance of a $\mathcal{Y}$-valued random variable $Y$ under $\mathcal{G}$, also known as sphericity, can then be formulated as ‘$\gamma(Y)$ is uniform on $S^{1}$’.
For $\mathcal{O}(2)$ or the general $d>2$ case, the group action is no longer free on each orbit. As a result there may be multiple group actions that carry $\iota\|y\|_{2}$ to a point $y$ on the hypersphere. While this may superficially seem like a potentially serious issue, we may simply view $\gamma(y)$ as uniformly drawn from all the ‘rotations’ that carry $\iota\|y\|_{2}$ to $y$. As a result, the only difference is that (10) will now hold almost surely, which suffices for our purposes.
A.2 Neyman–Pearson optimal e-values for $\mathbb{Q}$ on the sample space: the $t$-test and its generalizations
Suppose that $Y\sim\mathcal{N}_{d}(\mu\iota,I)$ on $\mathbb{R}^{d}\setminus\{0\}$, $\mu>0$ under the alternative and $Y$ is $\mathcal{G}$ invariant under the null. Let $\mathcal{G}$ be some compact group of orthonormal matrices; a subgroup of $\mathcal{O}(d)$.
Here, we conveniently have the Lebesgue measure as a $\mathcal{G}$ invariant reference measure, so that we may apply Proposition 2, which means we only need to consider the Gaussian density $q$ with respect to the Lebesgue measure when deriving optimal e-values:
$q(y)=1/(2\pi)^{d/2}\exp\left\{-\frac{1}{2}\|y-\iota\mu\|_{2}^{2}\right\}.$
By Theorem 6, the Neyman–Pearson optimal test rejects at level $\alpha$ when
$1/(2\pi)^{d/2}\exp\left\{-\frac{1}{2}\|y-\iota\mu\|_{2}^{2}\right\}$ $>q_{\alpha}^{\overline{G}}\left(1/(2\pi)^{d/2}\exp\left\{-\frac{1}{2}\|\overline{G}y-\iota\mu\|_{2}^{2}\right\}\right),$
where $\overline{G}$ is uniformly distributed on $\mathcal{G}$. This is equivalent to
$-y^{\prime}y+2\mu\iota^{\prime}y-\mu^{2}$ $>q_{\alpha}^{\overline{G}}\left(-y^{\prime}y+2\mu\iota^{\prime}\overline{G}y-\mu^{2}\right)$
so that the Neyman–Pearson optimal e-value / test may be concisely written as
$\varepsilon^{\textnormal{NP}}=1/\alpha\times\mathbb{I}\{\iota^{\prime}y$ $>q_{\alpha}^{\overline{G}}\left(\iota^{\prime}\overline{G}y\right)\},$ (11)
which is independent of $\mu$, so that this test is optimal against $\mathcal{N}(\iota\mu,I)$, uniformly in $\mu$.
Remark 13.
For $\mathcal{G}=\mathcal{O}(d)$, the test (11) is equal to the $t$-test by Theorem 6 in Koning and Hemerik [2023]. This matches the discussion in the final paragraphs of Lehmann and Stein [1949], who also conclude that the $t$-test is uniformly most powerful for testing spherical invariance against $\mathcal{N}_{d}(\mu\iota,I)$, $\mu>0$.
If $\mathcal{G}$ is a subgroup of $\mathcal{O}(d)$, this test may be viewed as a generalization of the $t$-test under weaker conditions; see Efron [1969] for an example in case of a group of sign-flips (diagonal matrices with diagonal elements in $\{-1,1\}$). Our results here show that the approach by Efron [1969] is most powerful for this sign-flipping group against Gaussianity.
Remark 14 (Optimality of the $t$-test beyond Lehmann and Stein [1949]).
If $\mathcal{G}=\mathcal{O}(d)$, then the $t$-test (11) may be reformulated as
$1/\alpha\times\mathbb{I}\{\iota^{\prime}y/\|y\|_{2}>q_{\alpha}^{\overline{G}}\left(\iota^{\prime}\overline{G}\iota\right)\},$
as $q_{\alpha}^{\overline{G}}\left(\iota^{\prime}\overline{G}y\right)=q_{\alpha}^{\overline{G}}\left(\iota^{\prime}\overline{G}\iota\|y\|_{2}\right)=\|y\|_{2}q_{\alpha}^{\overline{G}}\left(\iota^{\prime}\overline{G}\iota\right)$. Here, $\iota^{\prime}y/\|y\|_{2}$ may be interpreted as the correlation coefficient between $\iota$ and $y$.
Now, as the rejection event does not change if we apply a strictly increasing function to both sides, we may even conclude that the $t$-test is Neyman–Pearson-optimal for testing spherical invariance against any alternative with a density that is increasing in the correlation coefficient $\iota^{\prime}y/\|y\|_{2}$. This generalizes the result of Lehmann and Stein [1949], who only conclude optimality against Gaussian location shifts.
A.3 Log-optimal e-value
Following Corollary 2 and Proposition 2, the log-optimal e-value for $\mathcal{G}$-invariance against $\mathcal{N}(\iota\mu,I)$ is
$\varepsilon^{\textnormal{log}}(y)=\frac{q(y)}{\mathbb{E}_{\overline{G}}[q(\overline{G}y)]}=\frac{\exp\left\{\mu y^{\prime}\iota\right\}}{\mathbb{E}_{\overline{G}}\left[\exp\left\{\mu y^{\prime}\overline{G}\iota\right\}\right]}.$ (12)
While this may be viewed as the log-optimal version of the $t$-test, it is not uniformly log-optimal in $\mu$.
If $\mathcal{G}=\mathcal{O}(d)$, it is also not uniformly log-optimal in the class of alternatives with densities increasing in $\iota^{\prime}y/\|y\|_{2}$ as in Remark 14. Echoing Example 7, this underlines that there is no unique ‘e-value version’ of the $t$-test, nor even a unique ‘log-optimal’ version of the $t$-test: any e-value based on an alternative density $q$ that is non-decreasing in $\iota^{\prime}y/\|y\|_{2}$ may qualify. The underlying ‘problem’ is that the original $t$-test is Neyman–Pearson optimal uniformly against a large composite alternative, but specifying a log-optimal variant requires us to be much more specific about our alternative, because we cannot leverage the invariance of the e-value under monotone transformations of the test statistic as in Remark 14.
A.4 Alternative on orbits
We may apply the ideas in Section 4.4 to slightly enlarge the class of alternatives under which (12) is uniformly log-optimal by passing to the conditional distribution on each orbit. The conditional distribution of $Y\sim\mathcal{N}_{d}(\mu\iota,I)$ on each orbit is proportional to $\exp(\mu\iota^{\prime}y)$, where $y$ is on the orbit with radius $\|y\|_{2}$. For $\|y\|_{2}=1$, this is also known as the von Mises-Fisher distribution. The log-optimal e-value on each orbit $\varepsilon_{|O}^{\textnormal{log}}:O\mapsto[0,\infty]$ indeed corresponds to $\varepsilon^{\textnormal{log}}$:
$\varepsilon_{|O}^{\textnormal{log}}(y)=\frac{\exp\left\{\mu y^{\prime}\iota\right\}}{\mathbb{E}_{\overline{G}}\left[\exp\left\{\mu y^{\prime}\overline{G}\iota\right\}\right]}.$
As a consequence $\varepsilon^{\textnormal{log}}$ is log-optimal against any mixture over such conditional distributions on orbits.
A.5 Alternative on $\mathcal{G}$
In this section, we reduce ourselves to $d=2$ and $SO(2)$, so that the group action is free and the group will be easy to represent. Following Section A.1, we use a bijection between the unit circle $S^{1}$ and $SO(2)$ to more conveniently formulate the group using $S^{1}$.
As an alternative on the group, we consider the projected normal distribution $\mathcal{P}\mathcal{N}_{2}(\mu\iota,I)$. This arises as the pushforward of the Gaussian through the inversion kernel: if $Y\sim\mathcal{N}_{2}(\mu\iota,I)$, then $\gamma(Y)=Y/\|Y\|_{2}\sim\mathcal{P}\mathcal{N}_{2}(\mu\iota,I)$. Its density with respect to the uniform distribution on $S^{1}$ is
$\frac{\exp\{-\tfrac{1}{2}\mu^{2}\}}{2\pi}\left(1+\mu\iota^{\prime}v\frac{\Phi(\mu\iota^{\prime}v)}{\phi(\mu\iota^{\prime}v)}\right),$ (13)
where $v\in S^{1}$, $\Phi$ is the normal cdf and $\phi$ the pdf (Presnell et al., 1998; Watson, 1983). For $\mu=0$, this reduces to $1/(2\pi)$; the Haar-density. As a consequence, log-optimal e-value for testing the Haar measure against this projected normal distribution is simply the likelihood ratio between (13) and $1/(2\pi)$:
$\varepsilon_{S^{1}}^{\textnormal{log}}=\exp\{-\tfrac{1}{2}\mu^{2}\}\left(1+\mu\iota^{\prime}v\frac{\Phi(\mu\iota^{\prime}v)}{\phi(\mu\iota^{\prime}v)}\right).$
This may also be expressed as an e-value on $\mathcal{Y}$ by mapping through the inversion kernel:
$\varepsilon_{S^{1}}^{\textnormal{log}}(\gamma(y))$ $=\exp\{-\tfrac{1}{2}\mu^{2}\}\left(1+\mu\iota^{\prime}\gamma(y)\frac{\Phi(\mu\iota^{\prime}\gamma(y))}{\phi(\mu\iota^{\prime}\gamma(y))}\right)$
$=\exp\{-\tfrac{1}{2}\mu^{2}\}\left(1+\mu\iota^{\prime}y/\|y\|_{2}\frac{\Phi(\mu\iota^{\prime}y/\|y\|_{2})}{\phi(\mu\iota^{\prime}y/\|y\|_{2})}\right)$
which is an increasing function in $\iota^{\prime}y/\|y\|_{2}$ if $\mu>0$.
A.6 Permutations and softmax
The log-optimal e-value in (12) is strongly related to the softmax function. Indeed, if we choose $\overline{G}$ to be uniform on permutation matrices (which form a subgroup of the orthonormal matrices) and choose the unit vector $\iota=(1,0,\dots,0)$, then (12) becomes
$\frac{\exp\left\{\mu y_{1}\right\}}{\tfrac{1}{d}\sum_{i=1}^{d}\exp\left\{\mu y_{i}\right\}}.$ (14)
This is exactly the softmax function with ‘inverse temperature’ $\mu\geq 0$. Hence, the softmax function can be viewed as a likelihood ratio statistic for testing exchangeability (permutation invariance) against $\mathcal{N}((\mu,0,\dots,0),I)$.
Remark 15.
A related e-value appears in unpublished early manuscripts of Wang and Ramdas [2022] and Ignatiadis et al. [2023], who consider a ‘soft-rank’ e-value of the type $\varepsilon_{T}$ as in (4) with the choice of statistic
$T(y)=\frac{\exp(\kappa y_{1})-\exp(\kappa\min_{j}y_{j})}{\kappa},$ (15)
under exchangeability, for some inverse temperature $\kappa>0$.
Interestingly, this ‘soft-rank’ e-value for $\kappa=\mu$ is larger than the softmax e-value (14) if and only if the softmax e-value is larger than 1. In fact, the same holds if we replace $\exp(\kappa\min_{j}y_{i})$ by any positive constant $c$, and the relationship flips if $c$ is negative. For a positive constant $c$, we would therefore expect the ‘soft-rank’ e-value to be more volatile.
A.7 Testing sign-symmetry
[…]
$\exp\{\mu\iota^{\prime}y\}/\mathbb{E}_{\overline{G}}\exp\{\mu\iota^{\prime}\overline{G}y\}=2\exp\{\mu y\}/\left[\exp\{\mu y\}+\exp\{-\mu y\}\right],$
This can be generalized to $\mathcal{Y}=\mathbb{R}^{d}$ and $\mathcal{G}=\{-1,1\}^{d}$ and $\iota=d^{-1/2}(1,\dots,1)^{\prime}$. The log-optimal e-value becomes
$\exp\{d^{-1/2}\mu\iota^{\prime}y\}/\mathbb{E}_{\overline{g}}\exp\{d^{-1/2}\mu\overline{g}^{\prime}y\}=\prod_{i=1}^{d}\exp\{d^{-1/2}\mu y_{i}\}/\mathbb{E}_{\overline{g}_{i}}\exp\{d^{-1/2}\mu\overline{g}_{i}y_{i}\},$ (16)
where $\overline{g}$ is a $d$-vector of i.i.d. Bernoulli distributed random variables on $\{-1,1\}$ with probability .5.
Remark 16.
A related e-value can be derived from de la Peña [1999],
$\exp\{Z-Z^{2}/2\}.$
This object can be connected to our likelihood ratio, by simply normalizing it by $\mathbb{E}_{\overline{g}}[\exp\{\overline{g}Z-(\overline{g}Z)^{2}/2\}]$:
$\exp\{Z-Z^{2}/2\}/$ $\mathbb{E}_{\overline{g}}[\exp\{\overline{g}Z-(\overline{g}Z)^{2}/2\}]$
$=2\exp\{Z-Z^{2}/2\}/\left[\exp\{-Z-Z^{2}/2\}+\exp\{Z-Z^{2}/2\}\right]$
$=2\exp\{Z\}/\left[\exp\{-Z\}+\exp\{Z\}\right].$
This transformation makes the resulting e-value exact by Theorem 3, so that our e-value for sign-symmetry can be interpreted as an exact variant of the de la Peña [1999]-style e-value. This was also observed by Vovk and Wang [2024].
Moreover, Ramdas et al. [2022a] characterize the class of admissible e-processes for testing symmetry, and show that the e-process based on de la Peña [1999] is inadmissible. This inadmissibility is also visible in our simulations, where we find it is strongly dominated by ours.
Remark 17 (Relationship to Vovk and Wang [2024]).
Vovk and Wang [2024] also study the e-value (16). While they motivate this e-value from its reminiscence to the e-value in the Gaussian vs Gaussian setting, I show that it is in fact optimal for sign-symmetry a Gaussian location-shift. They also consider a particular sign-e-value, that relies on the number of positive signs. This may be viewed as mapping the data through the inversion kernel, and then constructing an e-value based on a particular statistic (the number of positive signs). A third e-value they consider relies on the number of ranks of observations with positive signs. This may be viewed as considering invariance under a group of both permutations and sign-flips, then mapping through the inversion kernel to the rank-sign combinations, and then deriving an e-value based on a particular statistic.
Appendix B Impoverishing filtrations
In the context of exchangeability, Example 14 recovers the result that no powerful test martingales exist. Instead of passing to an e-process, as discussed in Section 5 and Remark 11, Vovk [2021] considers moving to a less-informative, ‘impoverished’ filtration by passing to the ranks of the data. The practical implication is that we may no longer look at the full data, but only at the ranks. In exchange, it turns out that we may recover powerful martingales.
In this section, we show how the impoverishment of a filtration works in the more general context of group invariance, for general statistics and for statistics that mimic the role that the ranks play in exchangeability. This relies on two key ingredients: a subgroup $\mathcal{F}\subseteq\mathcal{G}$ and a statistic $H:\mathcal{X}\to\mathcal{Z}$ for which the subgroup induces a group action on its codomain. To induce a group action on the codomain, we require for each $x^{1},x^{2}\in\mathcal{X}$,
$H(x^{1})=H(x^{2})\implies H(Fx^{1})=H(Fx^{2}),\textnormal{ for all }F\in\mathcal{F}.$
Writing $z=H(x)$, the group action on the codomain is then defined as $Fz=H(Fx)$, $x\in\mathcal{X}$. This condition holds if and only if $H$ is equivariant under this group action: $FH(x)=H(Fx)$, for every $x\in\mathcal{X}$, $F\in\mathcal{F}$ [Eaton, 1989]. For this reason, we will refer to it as an equivariant statistic.
Proposition 6 captures the key idea: measuring evidence against $\mathcal{F}$ invariance of $Z:=H(X)$ also measures evidence against $\mathcal{G}$ invariance of $X$. As a consequence, we may apply all our methodology to testing $\mathcal{F}$ invariance of $Z$ and still obtain a valid e-value for $\mathcal{G}$ invariance of $X$.
Proposition 6.
Let $\mathcal{F}$ be a subgroup of $\mathcal{G}$ and let $H:\mathcal{X}\to\mathcal{Z}$ be $\mathcal{F}$-equivariant. Then, an e-value that is valid for $\mathcal{F}$ invariance of $H(X)$ is also valid for $\mathcal{G}$ invariance of $X$.
Proof.
$\mathcal{G}$ invariance of $X$ implies $\mathcal{F}$ invariance of $X$, as $\mathcal{F}$ is a subgroup of $\mathcal{G}$. Moreover, $\mathcal{F}$ invariance of $X$ implies $\mathcal{F}$ invariance of $H(X)$, as this group action is well-defined through the second assumption. Hence, the hypothesis of $\mathcal{F}$ invariance of $H(X)$ is at least as large as that of $\mathcal{G}$ invariance of $X$. Hence, an e-value that is valid for the former is also valid for the latter. ∎
To apply this in the sequential context described in Section 6.1, we must be careful to consider a sequence $(H_{n})_{n\geq 0}$ of statistics that are appropriately glued together with a projection on its codomain $\mathcal{Z}^{n-1}$: $H_{n-1}(x^{n-1})=\textnormal{proj}_{\mathcal{Z}^{n-1}}(H_{n}(x^{n}))$, and a nested sequence of subgroups $\mathcal{F}_{n}\subseteq\mathcal{G}_{n}$. In Example 16, we illustrate this approach by showing how we may reduce from exchangeability (continuing from Example 14) to within-batch exchangeability (continuing from Example 15) by selecting a particular equivariant statistic.
Example 16 (Reducing to within-batch-exchangeability).
Suppose $X^{n}=(Y_{1},\dots,Y_{n})$ and that $X^{n}$ is exchangeable. We now consider a statistic $H_{n}$ that effectively censors $X^{n}$ so that we only observe it in batches. Let $b_{1},b_{2},\dots$ denote the observation numbers at which a batch is completed, and $B_{n}$ the number of completed batches at time $n$. Then, we define the statistic equal to the most recently arrived batch $H_{n}(X^{n})=X^{b_{i}}$, and its codomain $\mathcal{Z}=\mathcal{X}^{b_{i}}$, for all $b_{i}\leq n<b_{i+1}$, $i<B_{n}$.
To induce a group action, we pass from the group $\mathfrak{P}_{n}$ of all permutations to its subgroup $\mathcal{F}_{n}=\mathfrak{P}^{1}\times\mathfrak{P}^{2}\times\cdots\times\mathfrak{P}^{B_{n}}\times I$, where $\mathfrak{P}^{i}$ permutes the observations within the $i$th batch of data, and $I$ acts as the identity on the yet-to-be-completed batch.
It remains to verify that this indeed induces a group action. This means we need to verify $H_{n}(x_{1}^{n})=H_{n}(x_{2}^{n})$ implies $H_{n}(Fx_{1}^{n})=H_{n}(Fx_{2}^{n})$ for all $F\in\mathcal{F}_{n}$ and $x_{1}^{n},x_{2}^{n}\in\mathcal{X}^{n}$. This is equivalent to checking whether $x_{1}^{b_{i}}=x_{2}^{b_{i}}$ implies $H_{n}(Fx_{1}^{n})=H_{n}(Fx_{2}^{n})$, where $b_{i}\leq n<b_{i+1}$, $i<B_{n}$. This is indeed satisfied, because $F$ only acts on the already completed batches.
B.1 Reduction to a single orbit
Recall from Section 5 that admissible e-processes for group invariance may be viewed as infimums over orbit-wise martingales. If there exists just a single orbit, then this infimum drops out so that admissible e-processes are martingales. While settings with just a single orbit may seem practically irrelevant, we can reduce to a single orbit by finding an appropriate subgroup $\mathcal{F}\subseteq\mathcal{G}$ and accompanying $\mathcal{F}$-equivariant statistic $H$ such that $H(X)$ has only a single orbit under $\mathcal{F}$.
While this approach may be applied to other statistics, we focus on a particularly attractive example of such a statistic: the unique inversion kernel $\gamma:\mathcal{X}\to\mathcal{G}$, which is an equivariant (possibly randomized) statistic [Kallenberg, 2017]. Such an inversion kernel maps to the group, and the group acting on itself trivially has a single orbit: the group itself. Moreover, $\mathcal{G}$ is a subgroup of itself, so that Proposition 6 applies, and we may measure evidence against $\mathcal{G}$ invariance of $X$ by measuring evidence against $\mathcal{G}$ invariance of $\gamma(X)$.
By passing through such a statistic, we are effectively observing a draw $\widetilde{G}:=\gamma(X)$ from the group itself. Recall from Section 4.5 that an e-value on the group, $\varepsilon:\mathcal{G}\to[0,\infty]$, is valid for $\mathcal{G}$ invariance if and only if
$\mathbb{E}_{\overline{G}}[\varepsilon(\overline{G})]\leq 1.$
Analogously, an e-process $(\varepsilon_{n})_{n\geq 0}$ is valid with respect to some filtration if and only if
$\mathbb{E}_{\overline{G}}[\varepsilon_{\tau}(\overline{G})]\leq 1,$
for every stopping time $\tau$ that is adapted to the same filtration.
As mentioned, any admissible e-process for such a simple hypothesis is a martingale Ramdas et al. [2022a]. This means such an admissible e-process may be induced as a Doob martingale in the style of Koning and van Meer [2025], as discussed in Proposition 4 for an arbitrary filtration $(\mathcal{I}_{n})_{n\geq 0}$ through
$\varepsilon_{n}=\mathbb{E}_{\overline{G}}[\varepsilon(\overline{G})\mid\mathcal{I}_{n}].$
Instead of such a backwards-induction of a martingale, we may forwards-construct a martingale by imposing some additional structure as in Section 6.1. In Section 6.1, we assumed that we are to observe an increasingly rich sequence of data $(X^{n})_{n\geq 0}$, $X^{n}=\textnormal{proj}_{\mathcal{X}^{n}}(X^{n+k})$, $n,k\geq 0$. Moreover, we introduced a nested sequence of groups $(\mathcal{G}_{n})_{n\geq 0}$, $\mathcal{G}_{n}\subseteq\mathcal{G}_{n+1}$. These two sequences were made compatible with each other by assuming that each projection map $\textnormal{proj}_{\mathcal{X}^{n}}$ is $\mathcal{G}_{n}$-equivariant. We now additionally assume that the orbit representatives are chosen in a compatible manner: $[\textnormal{proj}_{\mathcal{X}^{n}}(x^{n+k})]=\textnormal{proj}_{\mathcal{X}^{n}}([x^{n+k}])$. This makes the inversion kernels compatible with each other, as it implies [^9] Since $\gamma_{n}(x^{n})[x^{n}]\overset{a.s.}{=}x^{n}=\textnormal{proj}_{\mathcal{X}^{n}}(x^{n+k})\overset{a.s.}{=}\textnormal{proj}_{\mathcal{X}^{n}}(\gamma_{n+k}(x^{n+k})[x^{n+k}])$, which, if $\gamma_{n+k}(x^{n+k})\in\mathcal{G}_{n}$ equals $\gamma_{n+k}(x^{n+k})\textnormal{proj}_{\mathcal{X}^{n}}([x^{n+k}])=\gamma_{n+k}(x^{n+k})[\textnormal{proj}_{\mathcal{X}^{n}}(x^{n+k})]=\gamma_{n+k}(x^{n+k})[x^{n}]$.
$\gamma_{n+k}(x^{n+k})\in\mathcal{G}_{n}\implies\gamma_{n}(x^{n})[x^{n}]\overset{a.s.}{=}\gamma_{n+k}(x^{n+k})[x^{n}],$
so that $\gamma_{n+k}$ is an inversion kernel for $\mathcal{G}^{n}$ acting on $\mathcal{X}^{n}$. Assuming the group action is free, this implies $\gamma_{n}(x)=\gamma_{n+k}(x)$, for $x\in\mathcal{X}^{n}$, by the uniqueness of the inversion kernel. If the group action is not free, $\gamma_{n+k}$ may be viewed as a randomized statistic, which then shares the same unique distribution as $\gamma_{n}$ on $\mathcal{G}^{n}$.
We may now apply the machinery from Section 6.1 to derive a test martingale by constructing a conditional e-value for $\gamma_{n}(X^{n})$ conditional on $\gamma_{n-1}(X^{n-1})$. We illustrate this process in Example 17 and 18.
Example 17 (Exchangeability and ranks).
Suppose we have data $X^{n}=(Y_{1},Y_{2},\dots,Y_{n})$ and $X^{n}$ is exchangeable for each $n$; invariant under the group $\mathcal{G}_{n}$ of permutations. This means $\mathcal{G}_{n-1}$ is a subgroup of $\mathcal{G}_{n}$. Next we must select some orbit representative, which in the case of exchangeability comes down to selecting some canonical order of the elements. If the elements are real-valued, or admit some other natural ordering, then it makes sense to sort the elements accordingly and use this as the orbit representative, but any ordering suffices.
For example, suppose that $X^{n}=7314$, which we use as a shorthand for $Y_{1}=7,Y_{2}=3,Y_{3}=1,Y_{4}=4$. Suppose the orbit representative is selected as $1347$, then $\textnormal{Rank}_{n}(X^{n})=4213$. The ranks $4213$ may be interpreted as encoding the permutation that instructs how the elements of the orbit representative $1347$ must be permuted in order to recover $X^{n}=7314$: $4213$ states that the 4th element of $1347$ should be placed in the first position, the 2nd element of $1347$ in the second position, the 1st element in the third position and the 3rd element in the fourth position. That is, it encodes a permutation group action ‘$\times$’: $4213\times 1347=7314$. This means $\textnormal{Rank}_{n}(X^{n})\times[X^{n}]=X^{n}$, which is exactly the definition of an inversion kernel $\gamma_{n}=\textnormal{Rank}_{n}$.
Now, let us consider the distribution of $\gamma_{n}(X^{n})\mid\gamma_{n-1}(X^{n-1})$. Conditional on the ranking $\textnormal{Rank}_{n-1}(X^{n-1})=312$ of the first $(n-1)$ elements, we have under exchangeability that the ranking of $n$ elements $\textnormal{Rank}_{n}(X^{n})$ is uniform on $\{3124,4123,4132,4231\}$. Hence, constructing a conditional e-value is equivalent to constructing an e-value that is valid under a uniform distribution on this set.
In the context of rank-based testing of exchangeability, it is common to focus on the rank of the most recently arrived element. Working out the above for each possible conditioning, it is straightforward to show that this ‘last rank’ is uniform on $\{1,\dots,n\}$, independently of $\textnormal{Rank}_{n-1}(X^{n-1})$. Moreover, conditionally on $\textnormal{Rank}_{n-1}(X^{n-1})$, this last rank entirely determines $\textnormal{Rank}_{n}(X^{n})$. Hence, the last rank is in bijection with the distribution of $\textnormal{Rank}_{n}(X^{n})$ given $\textnormal{Rank}_{n-1}(X^{n-1})$, so that we may equivalently construct a conditional e-value by constructing an e-value that is valid for $\textnormal{Rank}_{n}(X^{n})\sim\textnormal{Unif}(\{1,\dots,n\})$. The above shows what underlies this last-rank result that is popularly used in conformal prediction, and how it generalizes to other settings.
Example 18 (Sequential sphericity).
We now move to the setting where $X^{n}=(Y_{1},\dots,Y_{n})$ is a spherical random $n$-vector in $\mathbb{R}^{n}$. That is, it is invariant under the orthogonal group $\mathcal{G}_{n}$ of $n\times n$ orthonormal matrices under matrix multiplication. Let us choose the statistic $H_{n}(X^{n})=X^{n}/\|X^{n}\|_{2}$, which maps from $\mathbb{R}^{n}$ to the unit sphere in $n$ dimensions. Note that, $H_{n}$ is equivariant: $H_{n}(GX^{n})=GX^{n}/\|GX^{n}\|_{2}=GX^{n}/\|X^{n}\|_{2}=GH_{n}(X^{n})$, for any $G\in\mathcal{G}_{n}$, so that it indeed induces a group action on the unit sphere. Under this group action, it has just a single orbit: the unit sphere itself. Furthermore, we have the required projection, as we may recover $H_{n-1}(X^{n-1})$ from $H_{n}(X^{n})$ by linearly projecting it onto $\mathbb{R}^{n-1}$ and then dividing by the norm of the resulting vector.
Under sphericity of $X^{n}$, we have that $H_{n}(X^{n})$ is spherical on the unit hypersphere. Conditional on $H_{n-1}(X^{n-1})$, $H_{n}(X^{n})$ is uniform on a semi-unit circle, of points whose first $(n-1)$ coordinates are in the direction of $H_{n-1}(X^{n-1})$. Hence, a conditional valid e-value is an e-value that is valid for a uniform distribution on this semi-circle. We may generalize this to instead observing, say, two coordinates at each point in time, so that this instead becomes a uniform distribution on a semi-sphere.
The standard $t$-test setting is recovered by considering the statistic $\iota_{n}^{\prime}H_{n}(X^{n})$ for the unit vector $\iota_{n}=(1,\dots,1)/\sqrt{n}$. Its distribution and conditional distribution is worked out in detail in Appendix A of Koning and van Meer [2025].
Appendix C Counterexample invariance through statistic
[…]
Appendix D Example: exchangeability
In this section, we discuss a highly concrete toy example of permutations on a small and finite sample space. While not as statistically interesting as the examples in Section A, it is more tangible as the group itself is finite and easy to understand.
D.1 Exchangeability on a finite sample space
Suppose our sample space $\mathcal{Y}$ consists of the vectors $[1,2,3]$, $[1,1,2]$ and all their permutations. As a group $\mathcal{G}$, we consider the permutations on 3 elements, which we will denote by $\{abc,acb,bac,bca,cab,cba\}$. For example, $bac$ represents the permutation that swaps the first two elements.
The orbits are then given by all permutations of $[1,2,3]$ and $[1,1,2]$
$O_{[1,1,2]}=\{[1,1,2],[1,2,1],[2,1,1]\},$
and
$O_{[1,2,3]}=\{[1,2,3],[1,3,2],[2,1,3],[2,3,1],[3,1,2],[3,2,1]\}.$
As $\mathcal{Y}$-valued orbit representatives, we pick the unique element in the orbit that is sorted in ascending order: $[1,1,2]$ and $[1,2,3]$.
For simplicity, let us restrict ourselves to $O_{[1,2,3]}$ first. On this orbit, the inversion kernel $\gamma$ is defined as the unique permutation that brings the element $[1,2,3]$ to $z\in O_{[1,2,3]}$. Moreover, on this orbit, the null hypothesis then states that $\gamma(Y)$ is uniform on the permutations, which in this case is equivalent to the hypothesis that $Y$ is uniform on $O_{[1,2,3]}$.
Now let us restrict ourselves to $O_{[1,1,2]}$. On this orbit, there are multiple permutations that may bring a given element back to $[1,1,2]$. For example, both $bac$, as well as the identity permutation $abc$ bring $[1,1,2]$ to itself, so that the group action is not free. More generally, any permutation that brings $[1,1,2]$ to $z\in O_{[1,1,2]}$, can be preceded by $bac$, and the result still brings $[1,1,2]$ to $z\in O_{[1,1,2]}$. Even more abstractly speaking, let $\mathcal{S}_{[y]}=\{G\in\mathcal{G}:G[y]=[y]\}$ be the stabilizer subgroup of $[y]$ (the subgroup that leaves $[y]$ unchanged). Then, if $G^{*}\in\mathcal{G}$ carries $[y]$ to $y$, so does any element in $G^{*}\mathcal{S}_{[y]}$.
To construct the inversion kernel $\gamma$ on $O_{[1,1,2]}$, let $\overline{S}_{[y]}$ denote a uniform distribution on $\{abc,bac\}$, which is well-defined as $\mathcal{S}_{[y]}$ is a compact subgroup and so admits a Haar probability measure (see Lemma 3). Moreover, let $G_{y}$ be an arbitrary permutation that carries $[y]$ to $y$, say $G_{[1,1,2]}=abc$, $G_{[1,2,1]}=acb$ and $G_{[2,1,1]}=cba$. Then, we define the inversion kernel as $\gamma(y)=G_{y}\overline{S}_{[y]}$. Concretely, this means that $\gamma([1,1,2])\sim\text{Unif}(abc,bac)$, $\gamma([1,2,1])\sim\text{Unif}(acb,bca)$ and $\gamma([2,1,1])\sim\text{Unif}(cba,cab)$. If $Y$ is indeed uniform on $O_{[1,1,2]}$, then $G_{Y}$ is uniform on $\{abc,acb,cba\}$ and so $\gamma(Y)$ is uniform on $\mathcal{G}$.
The definition of $\gamma$ on the sample space $\mathcal{Y}=O_{[1,2,3]}\cup O_{[1,1,2]}$ is obtained by combining the definitions on the two separate orbits.
Appendix E Omitted proofs
E.1 Proof of Theorem 1
Proof.
[…]
$\mathbb{E}_{Y}[\varepsilon(Y)]=\mathbb{E}_{[Y]}\mathbb{E}_{\overline{G}}[\varepsilon(\overline{G}[Y])]=\mathbb{E}_{[Y]}\mathbb{E}^{\textnormal{Unif}(O_{[Y]})}[\varepsilon]\leq\mathbb{E}_{[Y]}[1]=1.$
For the $\implies$’ direction, we assume $\mathbb{E}_{P}[\varepsilon]\leq 1$ for every $\mathcal{G}$ invariant probability $P$. Fix an orbit $O\in\mathcal{Y}/\mathcal{G}$. An example of a $\mathcal{G}$ invariant probability is $\textnormal{Unif}(O)$. Hence, $\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon]\leq 1$. As $O$ is arbitrary, this must hold for every $O$. ∎
E.2 Proof of Theorem 2
Proof.
[…]
$\varepsilon_{\alpha}(Z)$ $=\tfrac{1}{\alpha}\mathbb{I}\left\{T(Z)>q_{\alpha}^{\overline{G}}[T(\overline{G}Z)]\right\}+\frac{c([Z])}{\alpha}\mathbb{I}\left\{T(Z)=q_{\alpha}^{\overline{G}}[T(\overline{G}Z)]\right\}$
$=\tfrac{1}{\alpha}\mathbb{I}\left\{T(Z)>q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right\}+\frac{c([z])}{\alpha}\mathbb{I}\left\{T(Z)=q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right\}.$
[…]
$\tfrac{1}{\alpha}\mathbb{I}\left\{T(Z)>q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right\}+\frac{c([z])}{\alpha}\mathbb{I}\left\{T(Z)=q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right\}$
[…]
Now, this implies
$\mathbb{E}_{\overline{G}^{*}}\,\varepsilon_{\alpha}(\overline{G}Z)$ $=\mathbb{E}_{\overline{G}^{*}}\,\tfrac{1}{\alpha}\mathbb{I}\left\{T(\overline{G}^{*}[z])>q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right\}+\frac{c([z])}{\alpha}\mathbb{I}\left\{T(\overline{G}^{*}[z])=q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right\}$
$=\tfrac{1}{\alpha}\mathbb{P}_{\overline{G}^{*}}\left(T(\overline{G}^{*}[z])>q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right)+\frac{c([z])}{\alpha}\mathbb{P}_{\overline{G}^{*}}\left(T(\overline{G}^{*}[z])=q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right).$
Choosing
$c([z])=\frac{1-\mathbb{P}_{\overline{G}^{*}}\left(T(\overline{G}^{*}[z])>q_{\alpha}^{\overline{G}}\left[T(\overline{G}[z])\right]\right)}{\mathbb{P}_{\overline{G}^{*}}\left(T(\overline{G}^{*}[z])=q_{\alpha}^{\overline{G}}[T(\overline{G}[z])]\right)},$
yields $\mathbb{E}_{\overline{G}^{*}}\,\varepsilon_{\alpha}(\overline{G}Z)=1$, which proves the claim. ∎
E.3 Proof of Theorem 3
Proof.
[…]
$\mathbb{E}_{\overline{G}}[\varepsilon_{T}(\overline{G}y)]=\mathbb{E}_{\overline{G}_{1}}\left[\frac{T(\overline{G}_{1}y)}{\mathbb{E}_{\overline{G}_{2}}T(\overline{G}_{2}\overline{G}_{1}y)}\right]=\mathbb{E}_{\overline{G}_{1}}\left[\frac{T(\overline{G}_{1}y)}{\mathbb{E}_{\overline{G}_{2}}T(\overline{G}_{2}y)}\right]=\frac{\mathbb{E}_{\overline{G}_{1}}T(\overline{G}_{1}y)}{\mathbb{E}_{\overline{G}_{2}}T(\overline{G}_{2}y)}=1,$
and applying Theorem 1. For the $\implies$’ direction, assume $\varepsilon$ is some exact e-value for $\mathcal{G}$ invariance. Choose $T=\varepsilon$, so that
$\varepsilon_{\varepsilon}(y)=\frac{\varepsilon(y)}{\mathbb{E}_{\overline{G}}\varepsilon(\overline{G}y)}=\varepsilon(y),$
where the final equality follows from the fact that $\mathbb{E}_{\overline{G}}\varepsilon(\overline{G}y)=1$, by Theorem 1. ∎
E.4 Proof of Proposition 1
Proof.
Let $O\in\mathcal{Y}/\mathcal{G}$ be some arbitrary orbit. Let $Z$ be some random variable on this orbit satisfying $\mathbb{E}_{Z}[T(Z)]=\mathbb{E}_{\overline{G}}[T(\overline{G}[Z])]$.
[…]
$\mathbb{E}_{Z}[\varepsilon_{T}(Z)]=\mathbb{E}_{Z}\left[\frac{T(Z)}{\mathbb{E}_{\overline{G}}T(\overline{G}Z)}\right]=\mathbb{E}_{Z}\left[\frac{T(Z)}{\mathbb{E}_{\overline{G}}T(\overline{G}z)}\right]=\frac{\mathbb{E}_{Z}T(Z)}{\mathbb{E}_{\overline{G}}T(\overline{G}z)}$
for any arbitrary $z\in O$. Pick $z=[z]$, and apply the assumption to obtain
$\mathbb{E}_{Z}[\varepsilon_{T}(Z)]=\frac{\mathbb{E}_{Z}T(Z)}{\mathbb{E}_{Z}T(Z)}=1.$
Now, as $O$ was arbitrarily chosen, this holds for every orbit in $\mathcal{Y}/\mathcal{G}$. As a consequence, this holds for any mixture over orbits as well, so that $\varepsilon_{T}$ is indeed valid for group invariance. ∎
E.5 Proof of Theorem 4
Proof.
It suffices to show that $\mathbb{E}_{\overline{G}^{(1)},\dots,\overline{G}^{(M)}}\mathbb{E}_{\overline{G}}[\varepsilon_{T}^{M}(\overline{G}z)]=1$, for some $z$ on each orbit $O$. The strategy is to first show that the tuple $(T(\overline{G}z)$, $T(\overline{G}^{(1)}z)$, $\dots$, $T(\overline{G}^{(M)}z))$ is exchangeable, and to then apply Proposition 1.
Fix some arbitrary orbit $O\in\mathcal{Y}/\mathcal{G}$ and some element $z\in O$. Note that
$\varepsilon_{T}^{M}(\overline{G}z)=\frac{T(\overline{G}z)}{\frac{1}{M+1}\sum_{i=0}^{M}T(\overline{G}^{(i)}\overline{G}z)}.$
Central to this object is the tuple $(\overline{G},\overline{G}^{(1)}\overline{G},\dots,\overline{G}^{(M)}\overline{G})$.
Note that $(\overline{G}^{(1)}\overline{G},\dots,\overline{G}^{(M)}\overline{G})$ is independent from $\overline{G}$, as
[…]
As $(\overline{G}^{(1)},\dots,\overline{G}^{(M)})$ is mutually independent, we have that the tuple $(\overline{G},\overline{G}^{(1)}\overline{G},\dots,\overline{G}^{(M)}\overline{G})$ is mutually independent. Moreover, each element is marginally $\textnormal{Unif}(\mathcal{G})$, so that the tuple is i.i.d. and hence exchangeable.
We now prepare some notation to show that we may apply Theorem 1. Let us write $(\overline{G}_{*}^{(0)},\dots,\overline{G}_{*}^{(M)})=(\overline{G},\overline{G}^{(1)}\overline{G},\dots,\overline{G}^{(M)}\overline{G})$, so that
$\mathbb{E}_{\overline{G}^{(1)},\dots,\overline{G}^{(M)}}\left[\mathbb{E}_{\overline{G}}\left[\frac{T(\overline{G}z)}{\frac{1}{M+1}\sum_{i=0}^{M}T(\overline{G}^{(i)}\overline{G}z)}\right]\right]=\mathbb{E}_{\overline{G}_{*}^{(0)},\overline{G}_{*}^{(1)},\dots,\overline{G}_{*}^{(M)}}\left[\frac{T(\overline{G}_{*}^{(0)}z)}{\frac{1}{M+1}\sum_{i=0}^{M}T(\overline{G}_{*}^{(i)}z)}\right].$
Define the tuple
$\mathcal{T}:=(T(\overline{G}_{*}^{(0)}z),T(\overline{G}_{*}^{(1)}z),\dots,T(\overline{G}_{*}^{(M)}z)),$
which is exchangeable as $(\overline{G}_{*}^{(0)},\dots,\overline{G}_{*}^{(M)})$ is exchangeable. Moreover, define the statistic $S$ that returns the first element of such a tuple as $\mathcal{T}$. Finally, let $\overline{P}$ denote a uniform permutation on tuples of $(M+1)$ elements. Then, by Proposition 1,
$\mathbb{E}_{\overline{G}_{*}^{(0)},\overline{G}_{*}^{(1)},\dots,\overline{G}_{*}^{(M)}}\left[\frac{T(\overline{G}_{*}^{(0)}z)}{\frac{1}{M+1}\sum_{i=0}^{M}T(\overline{G}_{*}^{(i)}z)}\right]=\mathbb{E}_{\overline{P}}\left[\frac{S(\overline{P}\mathcal{T})}{\mathbb{E}_{\overline{P}}[S(\overline{P}\mathcal{T})]}\right]=1.$
E.6 Proof of Theorem 5
Proof.
The first claim follows from Theorem 1. For the second claim, suppose that $\varepsilon^{\prime}\in F_{+}^{\mathcal{Y}}$ is some other e-value that is valid for $\mathcal{G}$ invariance. Then, for each $O\in\mathcal{Y}/\mathcal{G}$, $\varepsilon^{\prime}_{|O}$ is also a valid e-value for $\mathrm{Unif}(O)$ by Theorem 1. Hence, by the assumption on $\varepsilon^{*}$,
$K_{O}\left(\varepsilon_{|O}^{*}\right)\geq K_{O}\left(\varepsilon^{\prime}_{|O}\right),\quad\text{for every }O\in\mathcal{Y}/\mathcal{G}.$
Since $\Psi$ is non-decreasing in each of its inputs, it follows that
$K(\varepsilon^{*})\equiv\Psi\left(\bigl{(}K_{O}\bigl{(}\varepsilon_{|O}^{*}\bigr{)}\bigr{)}_{O\in\mathcal{Y}/\mathcal{G}}\right)\geq\Psi\left(\bigl{(}K_{O}\bigl{(}\varepsilon^{\prime}_{|O}\bigr{)}\bigr{)}_{O\in\mathcal{Y}/\mathcal{G}}\right)\equiv K(\varepsilon^{\prime}),$
which proves the second claim. ∎
E.7 Proof of Theorem 6
Proof.
The strategy is to first use Theorem 5 with $K_{O_{[Y]}}(\cdot)=\mathbb{E}^{\mathbb{Q}}[\cdot\wedge 1/\alpha\mid O_{[Y]}]$ to decompose the problem into a problem on each orbit. Then, we apply the Neyman–Pearson lemma on each orbit, and finally we rewrite the result in terms of the densities $q$ and $\overline{q}$.
Note that both $\mathbb{E}^{\mathbb{Q}}$ and $\mathbb{E}^{\overline{\mathbb{Q}}}$ can be decomposed into a conditional expectation on the orbit $O_{[Y]}$ and a marginal distribution over the orbit $O_{[Y]}$ (or orbit representative $[Y]$):
$\mathbb{E}^{\mathbb{Q}}=\mathbb{E}_{[Y]}^{\mathbb{Q}}\left[\mathbb{E}^{\mathbb{Q}}[\,\cdot\mid O_{[Y]}]\right],$
$\mathbb{E}^{\overline{\mathbb{Q}}}=\mathbb{E}_{[Y]}^{\overline{\mathbb{Q}}}\left[\mathbb{E}^{\overline{\mathbb{Q}}}[\,\cdot\mid O_{[Y]}]\right].$
In fact, the marginal distribution over the orbit representative is the same under both $\overline{\mathbb{Q}}$ and $\mathbb{Q}$, by construction of $\overline{\mathbb{Q}}$. Let us fix a marginal density over the orbits $q^{\textnormal{orb}}$.
Given this density, we remove a set of null orbits $\mathcal{O}_{0}:=\{O\in\mathcal{Y}/\mathcal{G}:q^{\textnormal{orb}}(O)=0\}$. For the excluded points $y\in O\in\mathcal{O}_{0}$, we simply define $\varepsilon_{|O}^{\textnormal{NP}}(y)=1$, which is trivially an exact e-value for $\textnormal{Unif}(O)$ on such orbits.
For the non-null orbits $O\in\mathcal{O}_{+}:=(\mathcal{Y}/\mathcal{G})\setminus\mathcal{O}_{0}$, we have $q^{\textnormal{orb}}(O)>0$. Having fixed $q^{\textnormal{orb}}$, we may $[Y]$-uniquely determine the regular conditional expectations $\mathbb{E}^{\mathbb{Q}}[\,\cdot\mid O_{[Y]}]$ and $\mathbb{E}^{\overline{\mathbb{Q}}}[\,\cdot\mid O_{[Y]}]=\textnormal{Unif}(O_{[Y]})$. On these orbits, we use the conditional expectations to define densities $q^{O}=:q(\cdot\mid O)$ and $\overline{q}^{O}:=\overline{q}(\cdot\mid O)$. Expressed in such densities, a Neyman–Pearson optimal test/e-value at level $\alpha$ equals
$\varepsilon_{|O}^{\textnormal{NP}}(y)=\begin{cases}1,&\text{ if }q^{O}(y)=\overline{q}^{O}(y)=0,\\ 1/\alpha,&\text{ if }q^{O}(y)>c_{\alpha}^{O}\overline{q}^{O}(y),\\ k^{O},&\text{ if }q^{O}(y)=c_{\alpha}^{O}\overline{q}^{O}(y),\\ 0&\text{ if }q^{O}(y)<c_{\alpha}^{O}\overline{q}^{O}(y),\end{cases}$
for some constant $k^{O}$ and $c_{\alpha}^{O}$. This simply follows from the Neyman–Pearson lemma. By Theorem 5, we may now combine these restricted optimal e-values into an unrestricted optimal e-value.
Finally, recall that our goal is to restate this unrestricted optimal e-value in terms of the unconditional densities $\overline{q}$ and $q$, as
$\varepsilon^{\textnormal{NP}}(y)=\begin{cases}1,&\text{ if }q(y)=\overline{q}(y)=0,\\ 1/\alpha,&\text{ if }q(y)>c_{\alpha}^{O_{y}}\overline{q}(y),\\ k^{O_{y}},&\text{ if }q(y)=c_{\alpha}^{O_{y}}\overline{q}(y),\\ 0&\text{ if }q(y)<c_{\alpha}^{O_{y}}\overline{q}(y).\end{cases}$
First, note that on an omitted null orbit, $O\in\mathcal{O}_{0}$, we necessarily have $\overline{q}(y)=q(y)=0$, so that the fact that we imposed $\varepsilon^{\textnormal{NP}}(y)=1$ is properly captured by the $q(y)=\overline{q}(y)=0$ case. Next, on the non-omitted orbits, we have $q(y)=q^{O_{y}}(y)q^{\textnormal{orb}}(O_{y})$ and $\overline{q}(y)=\overline{q}^{O_{y}}(y)q^{\textnormal{orb}}(O_{y})$, and $q^{\textnormal{orb}}(O_{y})>0$, so that we may simply divide out $q^{\textnormal{orb}}(O_{y})$ to see that this equals the restricted e-value. ∎
E.8 Proof of Proposition 3
Proof.
First, note that $\overline{\mathbb{Q}}$ is a $\mathcal{G}$ invariant measure by construction. Next, we show that $d\mathbb{Q}/d\overline{\mathbb{Q}}$ is an e-value for $\mathcal{G}$ invariance.
For every event $A$, we have, by definition of $\overline{\mathbb{Q}}$ and applying Tonelli’s theorem,
$\mathbb{E}^{\mathbb{Q}}\left[1_{A}(Y)\right]$ $=\mathbb{E}^{\overline{\mathbb{Q}}}\left[1_{A}(Y)\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}(Y)\right]=\mathbb{E}_{\overline{G}}\left[\mathbb{E}^{\mathbb{Q}}\left[1_{A}(\overline{G}[Y])\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}(\overline{G}[Y])\right]\right]$
$=\mathbb{E}^{\mathbb{Q}}\left[\mathbb{E}_{\overline{G}}\left[1_{A}(\overline{G}[Y])\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}(\overline{G}[Y])\right]\right].$
We now specialize this to a $\mathcal{G}$ invariant event. For any $\mathcal{G}$ invariant event $B$, $1_{B}(\overline{G}[Y])=1_{B}([Y])$ and $1_{B}([Y])=1_{B}(Y)$, so that
$\mathbb{E}^{\mathbb{Q}}\left[1_{B}([Y])\right]=\mathbb{E}^{\mathbb{Q}}\left[1_{B}([Y])\mathbb{E}_{\overline{G}}\left[\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}(\overline{G}[Y])\right]\right].$
This implies
$\mathbb{E}_{\overline{G}}\left[\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}(\overline{G}[Y])\right]=1,$
$[Y]$-almost surely under $\mathbb{Q}$. Next, as the Radon-Nikodym derivative $d\mathbb{Q}/d\overline{\mathbb{Q}}$ is defined up to $\mathbb{Q}$-null sets, we choose it to equal 1 on appropriate null sets, so that
$\mathbb{E}_{\overline{G}}\left[\frac{d\mathbb{Q}}{d\overline{\mathbb{Q}}}(\overline{G}y)\right]=1,$
for all $y$. By Theorem 1, this means $d\mathbb{Q}/d\overline{\mathbb{Q}}$ is indeed an e-variable for $\mathcal{G}$ invariance.
As $d\mathbb{Q}/d\overline{\mathbb{Q}}$ is a Radon-Nikodym derivative between $\mathbb{Q}$ and some dominating element in the null hypothesis, Theorem 4.1 in Larsson et al. [2024] implies it is log-optimal. ∎
E.9 Proof of Theorem 8
Proof.
Let $X$ be some random variable on $\mathcal{X}$. Suppose we have $\sup_{\tau}\mathbb{E}_{\overline{G}}[\varepsilon_{\tau}(\overline{G}x)]\leq 1$, for every $x\in\mathcal{X}$. Then
$1\geq\mathbb{E}_{X}\sup_{\tau}\mathbb{E}_{\overline{G}}[\varepsilon_{\tau}(\overline{G}X)]\geq\sup_{\tau}\mathbb{E}_{X}\mathbb{E}_{\overline{G}}[\varepsilon_{\tau}(\overline{G}X)]=\sup_{\tau}\mathbb{E}_{Y}[\varepsilon_{\tau}(Y)],$
for $Y=\overline{G}X$. As any $\mathcal{G}$ invariant random variable $Y$ may be written as $Y=\overline{G}X$ for some $X$, $(\varepsilon_{n})_{n\geq 0}$ is an e-process.
Now, for the converse, fix some orbit $O\in\mathcal{X}/\mathcal{G}$. Suppose that $\sup_{\tau}\mathbb{E}_{X}[\varepsilon_{\tau}(X)]\leq 1$ for every $\mathcal{G}$ invariant random variable $X$. One particular such random variable is $\overline{G}x$, for $x\in O$. Hence, we have $\sup_{\tau}\mathbb{E}_{\overline{G}}[\varepsilon_{\tau}(\overline{G}x)]\leq 1$, for every $x\in O$. Now, as $O$ is arbitrarily given and the orbits partition $\mathcal{X}$, we have that this holds for every $x\in\mathcal{X}$. This finishes the proof of the claim. ∎
E.10 Proof of Theorem 9
Proof.
We have
$\sup_{\tau}\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{\tau}]$ $=\sup_{\tau}\mathbb{E}^{\textnormal{Unif}(O)}[\textnormal{ess\,inf}_{O\in\mathcal{X}/\mathcal{G}}\,\varepsilon_{\tau}^{O}]$
$=\sup_{\tau}\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{\tau}^{O}]\leq 1,$
where the first equality follows from the definition of $\varepsilon_{\tau}$, the second equality from the fact that $\varepsilon_{n}^{O}(x)=\infty$ for $x\not\in O$, and the inequality from Doob’s optional stopping theorem for non-negative supermartingales, and the assumption that the supermartingale starts at 1. ∎
E.11 Proof of Proposition 4
Proof.
As $\varepsilon^{*}$ is valid for $\mathcal{G}$ invariance, its orbital-restriction $\varepsilon_{|O}^{*}$ is valid for $\textnormal{Unif}(O)$. Moreover $\varepsilon_{n}^{O}$ is a martingale for $\textnormal{Unif}(O)$, as
$\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{n}^{O}\mid\mathcal{I}_{n-1}]=\mathbb{E}^{\textnormal{Unif}(O)}[\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{|O}^{*}\mid\mathcal{I}_{n}]\mid\mathcal{I}_{n-1}]=\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{|O}^{*}\mid\mathcal{I}_{n-1}]=\varepsilon_{n-1}^{O}.$
In addition, $\varepsilon_{0}^{O}=\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{|O}^{*}\mid\mathcal{I}_{0}]=\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{|O}^{*}]\leq 1$, as $\varepsilon^{*}$ is assumed to be valid for $\mathcal{G}$ invariance, which means $\varepsilon_{|O}^{*}$ is valid for $\textnormal{Unif}(O)$. Hence, by Theorem 9, its running infimum is an e-process for $\mathcal{G}$ invariance. Finally, if $\mathcal{I}_{N}=\mathcal{I}$, then $\varepsilon_{N}^{O}=\mathbb{E}^{\textnormal{Unif}(O)}[\varepsilon_{|O}^{*}\mid\mathcal{I}_{N}]=\varepsilon_{|O}^{*}$. Hence,
$\varepsilon_{N}=\textnormal{ess\,inf}_{O\in\mathcal{X}/\mathcal{G}}\,\varepsilon_{N}^{O}=\textnormal{ess\,inf}_{O\in\mathcal{X}/\mathcal{G}}\,\varepsilon_{|O}^{*}=\varepsilon^{*},$
by definition of $\varepsilon_{|O}^{*}$ as given in (8). ∎
E.12 Proof of Lemma 2
To prove Lemma 2, we prove a more general result in Proposition 7. Lemma 2 is recovered by choosing $h$ equal to the relevant projection map.
Let $\mathcal{Y}$ be our sample space on which our group $\mathcal{G}$ acts. Let $\mathcal{Z}$ be some other space. Suppose $h:\mathcal{Y}\to\mathcal{Z}$ is continuous. Assume $h$ induces a group action on $\mathcal{Z}$. That is, we assume $h(y_{1})=h(y_{2})\implies h(Gy_{1})=h(Gy_{2})$, for all $G\in\mathcal{G}$ and $y_{1},y_{2}\in\mathcal{Y}$. This means $h$ is equivariant for this group action on $\mathcal{Z}$: $h(Gy)=Gh(y)$.
Our goal is to characterize the conditional distribution of $Y\mid(O_{Y},h(Y))$ by a subgroup of $\mathcal{G}$. We start by characterizing the subgroup, and showing that it is compact.
Let us consider the subset $\mathcal{K}^{h}$ of $\mathcal{G}$ that stabilizes the statistic $h$ of the data:
$\mathcal{K}^{h}(y)=\{G\in\mathcal{G}:h(Gy)=h(y)\}=\{G\in\mathcal{G}:Gh(y)=h(y)\},$
Such a set $\mathcal{K}^{h}(y)$ is also known as a stabilizer subgroup. The fact that it is indeed a subgroup, and crucially its compactness are captured in Lemma 3.
Lemma 3.
$\mathcal{K}^{h}(y)$ is a compact subgroup of $\mathcal{G}$.
Proof.
We start by showing that $\mathcal{K}^{h}(y)$ is a subgroup. First, the identity $I$ is trivially in $\mathcal{K}^{h}(y)$. For any $K_{1},K_{2}\in\mathcal{K}^{h}(y)$, it is closed under composition: $K_{1}K_{2}h(y)=K_{1}h(y)=h(y)$. Moreover, for any $K\in\mathcal{K}^{h}(y)$, it contains its inverse $K^{-1}$: $h(y)=Ih(y)=K^{-1}Kh(y)=K^{-1}h(y)$.
Next, we show that $\mathcal{K}^{h}(y)$ is topologically closed. Define the map $f_{y}:\mathcal{G}\to\mathcal{Z}$ as the composition between $h$ and the group action: $f_{y}(G)=h(Gy)$. As both $h$ and the group action are continuous, their composition $f_{y}$ is also continuous. Since we latently assume any space we consider is Hausdorff, $\mathcal{Z}$ is also a $T_{1}$ space, so that $\{h(y)\}$ is closed. Hence, $\mathcal{K}^{h}(y)$ is the pre-image of the closed set $\{h(y)\}$ under a continuous map, and so $\mathcal{K}^{h}(y)$ is also closed. As $\mathcal{K}^{h}(y)$ is a closed subset of the compact set $\mathcal{G}$, it is also compact. ∎
In Proposition 7, we use this subgroup to characterize the conditional distribution $Y\mid(h(Y),O_{Y})$.
Proposition 7.
Let $Y$ be $\mathcal{G}$ invariant and $h:\mathcal{Y}\to\mathcal{Z}$ be $\mathcal{G}$ equivariant. For any orbit $O\in\mathcal{Y}/\mathcal{G}$ and $z\in h(O)$, pick $x\in O$ with $h(x)=z$. Let $\overline{K}^{h}\sim\textnormal{Unif}(\mathcal{K}^{h}(x))$, independent of $Y$. Then,
[…]
Proof.
We start by characterizing the orbit of $x$ under $\mathcal{K}^{h}(x)$,
$\mathcal{K}^{h}(x)x:=\{Kx:K\in\mathcal{K}^{h}(x)\}=\{Gx:G\in\mathcal{G},h(Gx)=z\}=\{y\in O:h(y)=z\}.$
[…]
Appendix F Full table hot hand application
Trigger 1 hit 2 hits 3 hits
$\beta$ $\beta$ $\beta$
Shooter ID 0.85 0.90 0.85 0.90 0.85 0.90
101 0.163 0.323 0.409 0.572 0.674 0.782
102 1.040 1.089 0.732 0.832 0.758 0.838
103 2.737 2.068 1.582 1.414 1.316 1.232
104 0.949 1.025 0.627 0.753 0.998 1.004
105 0.647 0.804 0.990 1.018 0.898 0.941
106 4.695 2.962 2.543 1.934 2.356 1.807
107 5.765 3.346 3.105 2.184 2.230 1.732
108 1.040 1.065 1.675 1.426 1.284 1.191
109 2.338 1.840 3.100 2.176 3.181 2.195
110 0.382 0.565 0.675 0.799 0.735 0.834
111 1.318 1.284 1.529 1.378 1.409 1.286
112 0.490 0.667 0.621 0.751 0.849 0.907
113 0.242 0.418 0.391 0.559 0.509 0.655
114 1.427 1.358 1.187 1.167 1.169 1.136
201 0.613 0.779 0.924 0.979 0.764 0.850
202 1.938 1.636 1.090 1.085 1.099 1.073
203 3.076 2.227 1.156 1.135 1.201 1.142
204 0.548 0.711 0.909 0.954 0.971 0.986
205 0.441 0.616 1.001 1.018 0.725 0.816
206 0.323 0.510 0.758 0.855 0.734 0.825
207 2.503 1.950 4.173 2.636 2.405 1.815
208 0.233 0.409 0.679 0.798 1.279 1.192
209 0.428 0.612 1.053 1.062 1.109 1.084
210 0.306 0.487 1.330 1.234 1.375 1.251
211 0.422 0.602 0.423 0.587 0.453 0.603
212 0.452 0.620 0.643 0.755 1.000 1.000
Product e-value 0.007 0.180 3.108 4.460 7.489 5.525
Table 2: Log-optimal e-values for each shooter in the controlled shooting experiment of Gilovich et al. [1985] for exchangeability against several hot hand alternatives, triggering after 1-3 hits for a modest effect $(\beta=0.85)$ and weak effect $(\beta=0.9)$. The final row reports the product e-value of the corresponding column.